### v2

# Automata Learning from Preference and Equivalence Queries

Abstract
Active automata learning from membership and equivalence queries is a foundational problem with numerous applications. We propose a novel variant of the active automata learning problem: actively learn finite automata using preference queries—i.e., queries about the relative position of two sequences in a total order—instead of membership queries. Our solution is Remap, a novel algorithm which leverages a symbolic observation table along with unification and constraint solving to navigate a space of symbolic hypotheses (each representing a set of automata), and uses satisfiability-solving to construct a concrete automaton (specifically a Moore machine) from a symbolic hypothesis. Remap is guaranteed to correctly infer the minimal automaton with polynomial query complexity under exact equivalence queries, and achieves PAC–identification ($\varepsilon$-approximate, with high probability) of the minimal automaton using sampling-based equivalence queries. Our empirical evaluations of Remap on the task of learning reward machines for two reinforcement learning domains indicate Remap scales to large automata and is effective at learning correct automata from consistent teachers, under both exact and sampling-based equivalence queries.
Introduction
Active automata learning has applications from software engineering (Schuts, Hooman, and Vaandrager 2016; Aarts et al. 2012) and verification (Lin et al. 2014) to interpretable machine learning (Weiss, Goldberg, and Yahav 2019) and learning reward machines (Tappler et al. 2019; Gaon and Brafman 2020; Xu et al. 2021; Dohmen et al. 2022). The classical problem formulation involves a teacher with access to a regular language and a learner which asks membership and equivalence queries to infer a finite automaton describing the regular language (Angluin 1987; Gold 1978).
Consider an alternative formulation: learning a finite automaton from preference and equivalence queries [^1] Shah et al. (2023) primarily investigates choosing between membership and preference queries. A preference query resolves the relative position of two sequences in a total order available to the teacher. Learning from preference data has applications in fine-tuning language models (Ouyang et al. 2022), learning conditional preference networks (Koriche and Zanuttini 2009; Guerin, Allen, and Goldsmith 2013), inferring reinforcement learning policies (Christiano et al. 2017), and learning Markovian reward functions (Bıyık, Talati, and Sadigh 2022; Sadigh et al. 2017; Bewley and Lécué 2021; Kalra and Brown 2023).
Figure 1: REMAP Algorithm Overview
Applications for learning finite automata from preferences over sequences include inferring sequence classifications (e.g. program executions (Giannakopoulou, Rakamaric, and Raman 2012), vehicle maneuvers, human-robot interactions) using ordered classes (e.g. safe, risky, dangerous, fatal), learning preference models (Herin, Perny, and Sokolovska 2023, 2024), distilling interpretable preference models (Weiss, Goldberg, and Yahav 2019), and inferring reward machines. However, no method currently exists for learning finite automata from preferences with the termination and correctness guarantees enjoyed by classical automata learning algorithms such as $\textrm{L}^{*}$ (Angluin 1987).
Unfortunately, adapting $\textrm{L}^{*}$ to the preference-based setting is challenging. Preference queries do not directly provide the concrete observations available from membership queries, as required by $\textrm{L}^{*}$, so our solution Remap addresses this challenge through a symbolic approach to $\textrm{L}^{*}$.
Remap (Figure 1) features termination and correctness guarantees for exact and probably approximately correct (Valiant 1984) (PAC) identification (Angluin 1988) of the desired automaton, with a strong learner capable of symbolic reasoning and constraint-solving to offset the weaker preference-based signals from the teacher. By using unification to navigate the symbolic space of hypotheses, and constraint-solving to construct a concrete automaton from a symbolic hypothesis, Remap identifies, in a polynomial number of queries, the minimal Moore machine isomorphic to one describing the teacher’s total order when using exact equivalence queries. Using sampling-based equivalence queries achieves PAC–identification. Our empirical evaluations apply Remap to learn reward machines for sequential-decision making domains in the reward machine literature. We measure query complexity for the exact and PAC–identification settings, and measure empirical correctness for the PAC–identification setting.
We contribute (a) Remap, a novel $\textrm{L}^{*}$ style algorithm for learning Moore machines using preference and equivalence queries; under exact and PAC-identification settings we provide (b) theoretical analysis of query complexity, correctness, and minimalism, and (c) supporting empirical results, demonstrating the efficacy of the algorithm’s ability to learn reward machines from preference and equivalence queries.
Background
Figure 2: Differences between Remap and $\textrm{L}^{*}$. Full example: Appendix, Figure 6. In (a) $\textrm{L}^{*}$ records concrete values from membership queries in its observation table, whereas (c,d,f) Remap uses SymbolicFill to perform preference queries and record variables in its symbolic observation table $\langle S,E,T;\mathcal{C},\Gamma\rangle$. Part (c) initializes $\langle S,E,T;\mathcal{C},\Gamma\rangle$; (d) expands $S$ with sequence $b$, followed by SymbolicFill, yielding a unified, closed, and consistent $\langle S,E,T;\mathcal{C},\Gamma\rangle$; (e) MakeHypothesis yields concrete hypothesis $h_{1}$ from the symbolic hypothesis and constraints in $\mathcal{C}$; (f) submits $h_{1}$ via equivalence query, receives and processes counterexample $bab$ with feedback $f(bab)=0$ from teacher by adding $bab$, $ba$, and $b$ to $S$, performs a SymbolicFill, sets the value of equivalence class for $T(bab)$ to $f(bab)$, yielding an inconsistent table. Part (b) shows the ground truth Moore machine with $\Sigma^{I}=\{a,b\},\Sigma^{O}=\{0,1\}$. Green symbolizes new information. Other colors in (f) visually highlight inconsistencies.
Prior to introducing Remap, we provide preliminaries of finite automata, followed by a discussion of Angluin’s $\textrm{L}^{*}$ algorithm in order to highlight how Remap differs from $\textrm{L}^{*}$.
Finite Automata  Automata describe sets of sequences. An alphabet $\Sigma$ is a set whose elements can be used to construct sequences; $\Sigma^{*}$ represents the set of sequences of any length created from elements of $\Sigma$. A sequence $s\in\Sigma^{*}$ has integer length $|s|\geq 0$; if $|s|=0$, then $s$ is the empty sequence $\varepsilon$. An element $\sigma\in\Sigma$ has length 1; if $s,t\in\Sigma^{*}$, then $s\cdot t$ represents $s$ concatenated with $t$, with length $|s|+|t|$. Different types of finite automata have different semantics. Deterministic automata feature a deterministic transition function $\delta$ defined over a set of states $Q$ and an input alphabet $\Sigma^{I}$. An output alphabet $\Sigma^{O}$ may be present to label states or transitions using a labeling function $L$. The tuple $\langle Q,q_{0},\Sigma^{I},\Sigma^{O},\delta,L\rangle$ is a Moore machine (Moore 1956), where $q_{0}\in Q$ is the initial state, $\delta:Q\times\Sigma^{I}\rightarrow Q$ describes transitions, and $L:Q\rightarrow\Sigma^{O}$ associates outputs with states. Extended, $\delta:Q\times(\Sigma^{I})^{*}\rightarrow Q$, where $\delta(q,\varepsilon)=q$ and $\delta(q,\sigma\cdot s)=\delta(\delta(q,\sigma),s)$. In Mealy machines (Mealy 1955), outputs are associated with transitions instead, so $L:Q\times\Sigma^{I}\rightarrow\Sigma^{O}$. Mealy and Moore machines are equivalent (Fleischner 1977) and can be converted between one another. Reward machines are an application of Mealy machines in reinforcement learning and are used to express a class of non-Markovian reward functions.
Active Automaton Learning  Consider the problem of actively learning a Moore machine $\langle Q,q_{0},\Sigma^{I},\Sigma^{O},\delta,L\rangle$ to exactly model a function $f:(\Sigma^{I})^{*}\rightarrow\Sigma^{O}$, where $\Sigma^{I}$ and $\Sigma^{O}$ are input and output alphabets of finite size known to both teacher and learner. We desire a learner which learns a model $\hat{f}$ of $f$ exactly; that is for all $s\in(\Sigma^{I})^{*}$, we require $\hat{f}(s)=f(s)$, where $\hat{f}(s)=L(\delta(q_{0},s))$, with the assistance of a teacher $\mathcal{T}$ that can answer questions about $f$.
In Angluin’s seminal active learning algorithm, $\textrm{L}^{*}$ (Angluin 1987), the learner learns $\hat{f}$ as a binary classifier, where $|\Sigma^{O}|=2$, to determine sequence membership of a regular language by querying $\mathcal{T}$ with: (i) membership queries, where the learner asks $\mathcal{T}$ for the value of $f(s)$ for a particular sequence $s$, and (ii) equivalence queries, where the learner asks $\mathcal{T}$ to evaluate whether for all $s\in(\Sigma^{I})^{*}$, $\hat{f}(s)=f(s)$. For the latter query, $\mathcal{T}$ returns True if the statement holds; otherwise a counterexample $c$ for which $\hat{f}(c)\neq f(c)$ is returned. An observation table $\langle S,E,T\rangle$ records the concrete observations acquired by the learner’s queries (see Figure 2a). Here, $S$ is a set of prefixes, $E$ is a set of suffixes, and $T$ is the empirical observation function that maps sequences to output values—$T:(S\cup(S\cdot\Sigma^{I}))\cdot E\rightarrow\Sigma^{O}$. The observation table is a two-dimensional array; $s\in(S\cup(S\cdot\Sigma^{I}))$ indexes rows, and $e\in E$ indexes columns, with entries given by $T(s\cdot e)$. Any proposed hypothesis $\hat{f}$ must be consistent with $T$. The algorithm operates by construction: a deterministic transition function must be found exhibiting consistency (deterministic transitions) and operates in a closed manner over the set of states. If the consistency or closure requirements are violated, then membership queries are executed to expand the observation table. Once a suitable transition function is found, a hypothesis $\hat{f}$ can be made and checked via the equivalence query. The algorithm terminates if $\hat{f}(s)=f(s)$ for all $s\in(\Sigma^{I})^{*}$; otherwise $\textrm{L}^{*}$ continues on by adding counterexample $c$ and all its prefixes to the table, and finds another transition function satisfying the consistency and closure requirements.
Consequently, we consider how $\textrm{L}^{*}$-style learning can be used to learn Moore machines from preference queries over sequences, as a foray into understanding how the structure of finite automata can be learned from comparison information.
Problem Statement
We consider how a Moore machine $\langle Q,q_{0},\Sigma^{I},\Sigma^{O},\delta,L\rangle$ can be actively learned from preference queries over $(\Sigma^{I})^{*}$. We focus on the case of a finite sized $\Sigma^{I}$ and $\Sigma^{O}$ known to both teacher and learner. With a preference query, the learner asks the teacher $\mathcal{T}$ which of two sequences $s_{1}$ and $s_{2}$ is preferred, or if both are equally preferable. This requires a preference model, which we assume represents a total ordering over $(\Sigma^{I})^{*}$; we also assume $\Sigma^{O}$ is totally ordered. Thus, we consider a preference model for $\mathcal{T}$ where $f:(\Sigma^{I})^{*}\rightarrow\Sigma^{O}$ is consistent with both orderings, i.e., $\mathcal{T}$ prefers $s_{1}$ over $s_{2}$ if $f(s_{1})>f(s_{2})$, or otherwise has equal preference if $f(s_{1})=f(s_{2})$.
Under this problem formulation, several options for evaluating hypothesis equivalence (is $\hat{f}\equiv f$?) can be defined. The standard option requires exact equivalence: $\hat{f}(s)=f(s)$ for all $s$ in $(\Sigma^{I})^{*}$. An alternative requires the total ordering implied by $\hat{f}$ and $f$ to match. That is, for all $s,t$ in $(\Sigma^{I})^{*}$, both $f(s)\mathbf{R}_{s,t}f(t)$ and $\hat{f}(s)\mathbf{R}_{s,t}\hat{f}(t)$ are satisfied, where $\mathbf{R}_{s,t}\in\{=,>,<\}$. Since a hypothesis $\hat{f}$ satisfying the standard definition always satisfies the alternative, we focus on exact equivalence queries where $\mathcal{T}$ returns feedback along with counterexample $c$. Feedback strength directly impacts how many hypotheses per counterexample can be eliminated by the learner: returning $\hat{f}(c)\neq f(c)$ is weak feedback compared to returning $\hat{f}(c)=f(c)$ or returning $\hat{f}(c)$ is in a subset of $\Sigma^{O}$. Although Remap outputs a concrete Moore machine, Remap navigates symbolic Moore machine222i.e., a Moore machine with symbolic values as outputs. space using solely preference information, while concrete information assists in selecting the concrete hypothesis. Thus, our theoretical analysis considers equivalence queries that provide counterexamples with strong feedback.
The Remap Algorithm
Remap is a $\textrm{L}^{*}$-based algorithm employing preference and equivalence queries to gather constraints. By first leveraging unification to navigate the symbolic hypothesis space of Moore machines, solving the constraints yields a concrete Moore machine. In particular, Remap (Algorithm 1) learns a Moore machine $\langle\hat{Q},\hat{q}_{0},\Sigma^{I},\Sigma^{O},\hat{\delta},\hat{L}\rangle$, representing a multiclass classifier $\hat{f}(s)=\hat{L}(\hat{\delta}(\hat{q}_{0},s))$.
Central to Remap is a new construct called a symbolic observation table, shown in Figure 2c, 2d, and 2f, as well as an associated algorithm for unification (Robinson 1965) inspired by Martelli and Montanari (1976) to contain the fresh variable explosion; both enable the learner to generate symbolic hypotheses purely from observed symbolic constraints, along with a constraint solver for obtaining a concrete hypothesis. We first discuss these core components of Remap and how they fit together, then illustrate the correctness and termination guarantees.
Symbolic Observation Tables
Recall that in classic $\textrm{L}^{*}$, the observation table entries are concrete values obtained from membership queries (Figure 2a). However, observations obtained from preference queries are constraints, rather than concrete values, indicating that for a pair of sequences $s_{1}$ and $s_{2}$, one of $f(s_{1})>f(s_{2})$, $f(s_{1})<f(s_{2})$, or $f(s_{1})=f(s_{2})$ holds. We therefore introduce the symbolic observation table $\langle S,E,T;\mathcal{C},\Gamma\rangle$, where $S$ is a set of prefixes, $E$ is a set of suffixes, $\mathcal{C}$ is the set of known constraints, $\Gamma$ is a context which uniquely maps sequences to variables, with the set of variables $\mathcal{V}$ in the context $\Gamma$ given by $\mathcal{V}=\{\Gamma[s\cdot e]|s\cdot e\in(S\cup(S\cdot\Sigma^{I}))\cdot E\}$, and $T:(S\cup(S\cdot\Sigma^{I}))\cdot E\rightarrow\Gamma$ maps queried sequences to variables. Thus, in a symbolic observation table, the entry for each prefix-suffix pair is a variable, rather than a concrete value; and the constraints over those variables are stored in $\mathcal{C}$ (Figure 2c, 2d, 2f). The constraints from the preferences of $\mathcal{T}$ over $f$ about $s_{1}$ and $s_{2}$ correspond with $T(s_{1})>T(s_{2})$, $T(s_{1})<T(s_{2})$, or $T(s_{1})=T(s_{2})$, respectively.
Definition 1.
An equivalence class $\mathbb{C}$ of variables is a set with the property that all members of $\mathbb{C}$ are equivalent to each other. The representative .$\mathbb{C}.\text{{rep}}$ of an equivalence class $\mathbb{C}$ is a deterministically elected member of $\mathbb{C}$. The set of variables $\mathcal{V}$ can be partitioned into disjoint equivalence classes. The set of equivalence classes $\mathcal{C}_{EC}$ corresponds with the partitioning of $\mathcal{V}$ into the smallest possible number of equivalence classes consistent with equality constraints. Let ={.$\mathcal{R}=\{\mathbb{C}.\text{{rep}}|\mathbb{C}\in\mathcal{C}_{EC}\}$ be the set of representatives.
Unification and Constraints
Since preference queries return observations comparing the values of two sequences, we leverage a simple unification algorithm to ensure the number of unique variables in the table remains bounded by $|\Sigma^{O}|$. Whenever we observe the constraint $f(s_{1})=f(s_{2})$, this implies $T(s_{1})=T(s_{2})$, so we add $\Gamma[s_{1}]$ and $\Gamma[s_{2}]$ to an equivalence class $\mathbb{C}$, and elect a representative from $\mathbb{C}$.
If $\Gamma[s_{1}]$ already belongs to an equivalence class $\mathbb{C}\in\mathcal{C}_{EC}$, but $\Gamma[s_{2}]$ is an orphan variable—belonging to no class in $\mathcal{C}_{EC}$—then $\Gamma[s_{2}]$ is merged into $\mathbb{C}$. Swap $s_{1}$ and $s_{2}$ for the other case. If $\Gamma[s_{1}]$ and $\Gamma[s_{2}]$ belong to separate classes $\mathbb{C}_{1}$ and $\mathbb{C}_{2}$, with $\mathbb{C}_{1}\neq\mathbb{C}_{2}$, then $\mathbb{C}_{1}$ and $\mathbb{C}_{2}$ are merged into one via $\mathbb{C}\leftarrow\mathbb{C}_{1}\cup\mathbb{C}_{2}$, and one of $\mathbb{C}_{1}.\text{{rep}}$ and $\mathbb{C}_{2}.\text{{rep}}$ is deterministically elected as the representative .$\mathbb{C}.\text{{rep}}$.
When this unification process is applied to a symbolic observation table $\langle S,E,T;\mathcal{C},\Gamma\rangle$, we are unifying $\mathcal{V}$ (the set of variables in the context $\Gamma$) according to the set of known equivalence constraints $\mathcal{C}_{EQ}$ and equivalence classes $\mathcal{C}_{EC}$ available in $\mathcal{C}$; and each variable in the table is replaced with its equivalence class representative (see Figure 2cdf after the large right curly brace). That is, for all $s$ in $(S\cup S\cdot\Sigma^{I})\cdot E$, we substitute $T(s)\leftarrow\mathbb{C}.\text{{rep}}$ if $T(s)\in\mathbb{C}$. In the resulting unified symbolic observation table, $T$ maps queried sequences to the set of representatives $\mathcal{R}$. Note $|\mathcal{R}|=|\mathcal{C}_{EC}|\leq|\Sigma^{O}|$.
Besides constraints from preferences queries, the learner obtains constraints about the value of $T(c)$ from equivalence queries (Figure 2e to 2f) and adds them to $\mathcal{C}$. When first obtained, these constraints may possibly be expressed in terms of orphan variables, but during the process of unification, each orphan variable joins an equivalence class and is then replaced by its equivalence class representative in the constraint. Thus, after unification, all constraints in $\mathcal{C}$ are expressed in terms of equivalence class representatives.
Finally, unifying the symbolic observation table is critical for making a symbolic hypothesis without knowledge of concrete value assignments: unification permits $\langle S,E,T;\mathcal{C},\Gamma\rangle$ to become closed and consistent—prerequisites for generating a symbolic hypothesis.
Definition 2.
A symbolic observation table is closed if $\textbf{{rows}}(S\cdot\Sigma^{I})\subseteq\textbf{{rows}}(S)$, where $\textbf{{rows}}(S)=\{\textbf{{row}}(s)|s\in S\}$, and $\textbf{{row}}(s)$ is the row in $\langle S,E,T;\mathcal{C},\Gamma\rangle$ indexed by $s$.
Definition 3.
A symbolic observation table is consistent if for all sequence pairs $s_{1}$ and $s_{2}$ where $\textbf{{row}}(s_{1})\equiv\textbf{{row}}(s_{2})$, all their transitions also remain equivalent with each other: $\textbf{{row}}(s_{1}\cdot\sigma)\equiv\textbf{{row}}(s_{2}\cdot\sigma)$ for all $\sigma\in\Sigma^{I}$.
Definition 4.
A symbolic observation table is unified if for all $s\cdot e\in(S\cup(S\cdot\Sigma^{I}))\cdot E$ it holds that $T(s\cdot e)\in\mathcal{R}$.
Making a Symbolic Hypothesis
If a symbolic observation table $\langle S,E,T;\mathcal{C},\Gamma\rangle$ is unified, closed, and consistent, a symbolic hypothesis can be made (Figure 2e). This construction is identical to $\textrm{L}^{*}$, except that the outputs are symbolic:
$\hat{Q}$ =$=\{\textbf{{row}}(s)|\forall s\in S\}\text{ \ is the set of states}$
$\hat{q}_{0}$ $=\textbf{{row}}(\varepsilon)\text{ \ is the initial state}$
$\hat{\delta}(\textbf{{row}}(s),\sigma)$ $=\textbf{{row}}(s\cdot\sigma)\text{ \ for all $s\in S$ and $\sigma\in\Sigma^{I}$}$
$\hat{L}(\textbf{{row}}(s))$ $=T(s\cdot\varepsilon)\text{ \ is the sequence to output function}$
$\langle\hat{Q},\hat{q}_{0},\Sigma^{I},\Sigma^{O},\hat{\delta},\hat{L}\rangle\text{ is a symbolic hypothesis.}$.
Making a Concrete Hypothesis
The learner finds a satisfying solution $\Lambda$ to the set of constraints $\mathcal{C}$, while subject to the global constraint requiring the value of each representative to be in $\Sigma^{O}$ (Figure 2e). Thus, $\hat{L}$ becomes concrete.
$\Lambda\longleftarrow\textsc{FindSolution}(\langle S,E,T;\mathcal{C},\Gamma\rangle,\Sigma^{O})$ (1)
$\hat{L}(\textbf{{row}}(s))=\Lambda[T(s\cdot\varepsilon)]$ (2)
In particular, $\Lambda$ finds satisfying values for each member of $\mathcal{R}$, the set of equivalence class representatives. Since $\langle S,E,T;\mathcal{C},\Gamma\rangle$ is unified, we are guaranteed that $T$ maps from queried sequences to $\mathcal{R}$; hence $\Lambda[T(s\cdot\varepsilon)]$ is guaranteed to resolve to a concrete value as long as the teacher provides consistent preferences.
Algorithm 1 Remap
Input: Alphabets $\Sigma^{I}$ (input) and $\Sigma^{O}$ (output), teacher $\mathcal{T}$
Output: Moore Machine $\mathcal{H}=\langle\hat{Q},\Sigma^{I},\Sigma^{O},\hat{q}_{0},\hat{\delta},\hat{L}\rangle$
1:  Initialize $\mathcal{O}=\langle S,E,T;\mathcal{C},\Gamma\rangle$ with $S=\{\varepsilon\},E=\{\varepsilon\}$, $\mathcal{C}=\{\}$, $\Gamma=\emptyset$
2:  $\mathcal{O}\longleftarrow$ SymbolicFill$\left(\mathcal{O}|\mathcal{T}\right)$
3:  repeat
4:     $\mathcal{O}\longleftarrow$ MakeClosedAndConsistent$\left(\mathcal{O}|\mathcal{T}\right)$
5:     $\mathcal{H}\longleftarrow$MakeHypothesis$\left(\mathcal{O},\Sigma^{I},\Sigma^{O}\right)$
6:     result $\longleftarrow$ EquivalenceQuery$(\mathcal{H}|\mathcal{T})$
7:     $\mathcal{O}\longleftarrow$ProcessCex$(\mathcal{O},\text{{result}})$ if result $\neq$ correct
8:  until result $=$ correct
9:  return  $\mathcal{H}$
Algorithm 2 A Query Efficient Symbolic Fill Procedure
procedure SymbolicFill$\left(\langle S,E,T;\mathcal{C},\Gamma\rangle|\mathcal{T}\right)$
[…]
2:  seqs $=$ PopulateMissingFreshVars($\mathcal{O}$)
3:  $\mathcal{O}$longleftarrow PrefQsByRandomizedQuicksortFollowedByLinearMerge(seqs, oldsortedseqs, $\mathcal{O}|\mathcal{T}$)
4:  $\mathcal{O}\longleftarrow$ Unification$(\mathcal{O})$
5:  return  $\mathcal{O}$
Remap
We now describe Remap (Algorithm 1) in terms of the previously discussed components. In order to make a symbolic hypothesis, Remap must first obtain a unified, closed, and consistent $\langle S,E,T;\mathcal{C},\Gamma\rangle$. To perform closed and consistency checks, the table must be unified. Therefore, Remap must symbolically fill $\langle S,E,T;\mathcal{C},\Gamma\rangle$ by asking preference queries and performing unification to obtain a unified table. If the unified table is not closed or not consistent, then the table is alternately expanded and symbolically filled until the table becomes unified, closed, and consistent.
SymbolicFill  A symbolic fill (Algorithm 2 and Appendix Algorithm 4 and 5) (1) creates fresh variables for empty entries in the table, (2) asks preference queries, and (3) performs unification. If a sequence $s\cdot e\in(S\cup(S\cdot\Sigma^{I}))\cdot\,E$ does not have an associated variable in the context $\Gamma$, then a fresh variable $\Gamma[s\cdot e]$ is created. Preference queries are executed to obtain the total ordering of $(S\cup(S\cdot\Sigma^{I}))\cdot E$. In our implementation, we use preference queries in place of comparisons in randomized quicksort and linear merge. Once every sequence in $(S\cup(S\cdot\Sigma^{I}))\cdot E$ has a variable, and once the preference queries have been completed, unification is performed. SymbolicFill is called on lines 2, 4 (in MakeClosedAndConsistent), and 7 (in ProcessCex.). Unification is shown in Appendix Algorithm 9.
Ensuring Consistency   If the unified table is not consistent, then there exists a pair $s_{1},s_{2}\in S$ and $\sigma\in\Sigma^{I}$ for which $\textbf{{row}}(s_{1})\equiv\textbf{{row}}(s_{2})$ and $\textbf{{row}}(s_{1}\cdot\sigma)\not\equiv\textbf{{row}}(s_{2}\cdot\sigma)$, implying there is an $e\in E$ such that $T(s_{1}\cdot\sigma\cdot e)\not\equiv T(s_{2}\cdot\sigma\cdot e)$. To attempt to make the table consistent, add $\sigma\cdot e$ to $E$, and then perform a symbolic fill. Figure 2f shows inconsistency.
Ensuring Closedness   If the unified table is not closed, then $\textbf{{rows}}(S\cdot\Sigma^{I})\not\subseteq\textbf{{rows}}(S)$. To attempt to make the table closed, find a row $\textbf{{row}}(s^{\prime})$ in $\textbf{{rows}}(S\cdot\Sigma^{I})$ but not in $\textbf{{rows}}(S)$. Add $s^{\prime}$ to $S$, update $S\cdot\Sigma^{I}$, then fill symbolically. Figure 2c to 2d shows a closure process.
The closed and consistency checks occur in a loop (consistency first, closed second) inside MakeClosedAndConsistent until the table becomes unified, closed, and consistent. Then hypothesis $h=\langle\hat{Q},\hat{q}_{0},\Sigma^{I},\Sigma^{O},\hat{\delta},\hat{L}\rangle$ is generated by MakeHypothesis (Figure 2e) and is sent to the teacher via EquivalenceQuery (Figure 2e to 2f). If $h$ is wrong, then a counterexample $c$ is returned, as well as a new constraint on the value of $\hat{f}(c)$. The counterexample $c$ and all its prefixes are added to $S$, and then a symbolic fill is performed. Then the constraint on the value of $\hat{f}(c)$ is added to $\mathcal{C}$ as a constraint on the value of the representative at $T(c)$.
Theoretical Guarantees of Remap
We now cover the algorithmic guarantees of Remap when $\mathcal{T}$ uses exact equivalence queries, and show how sampling-based equivalence queries achieves PAC–identification.
We first detail how Remap guarantees termination and yields a correct, minimal Moore machine that classifies sequence equivalently to $f$. If Remap terminates, then the final hypothesis must be correct, since termination occurs only if no counterexamples exist for the final hypothesis. Therefore, if the hypothesized Moore machine classifies all sequences correctly according to the teacher, it must be correct. Thus, proving termination implies correctness. See Appendix for sketches and proofs. Here, we assume the teacher provides feedback $\hat{f}(c)=f(c)$ with counterexample $c$.
Theorem 1.
If $\langle S,E,T;\mathcal{C},\Gamma\rangle$ is unified, closed, and consistent, and the range of MakeHypothesis$(\langle S,E,T;\mathcal{C},\Gamma\rangle)$ is $\mathcal{H}$, then every hypothesis $h\in\mathcal{H}$ is consistent with constraints $\mathcal{C}$. Any other hypothesis consistent with $\mathcal{C}$, but not contained in $\mathcal{H}$, must have more states.
Theorem 1 establishes that the output of Remap will be the smallest Moore machine consistent with all the constraints in $\mathcal{C}$. This is necessary to prove termination.
Lemma 1.
Whenever a counterexample $c$ is processed, either 00 or $1$ additional representative values becomes known.
Theorem 2.
Suppose $\langle S,E,T;\mathcal{C},\Gamma\rangle$ is unified, closed, and consistent. Let $\hat{h}=\textsc{MakeHypothesis}(\langle S,E,T;\mathcal{C},\Gamma\rangle)$ be the hypothesis induced by $\Lambda$, a satisfying solution to $\mathcal{C}$. If the teacher returns a counterexample $c$ as the result of an equivalence query on $\hat{h}$, then at least one of the following is true about $\hat{h}$: (a) $\hat{h}$ contains too few states, or (b) the satisfying solution $\Lambda$ inducing $\hat{h}$ is either incomplete or incorrect.
Corollary 1.
Remap must terminate when the number of states and number of known representative values in a concrete hypothesis reach their respective upper bounds.
Theorem 3 (Query Complexity).
If $n$ is the number of states of the minimal automaton isomorphic to the target automaton, and $m$ is the maximum length of any counterexample sequence that the teacher returns, then (a) Remap executes at most $n+|\Sigma^{O}|-1$ equivalence queries, and (b) the preference query complexity is $\mathcal{O}(mn^{2}\ln(mn^{2}))$, which is polynomial in the number of unique sequences queried.
Lemma 1, Theorem 2, and Corollary 1 imply Remap makes progress towards termination with every hypothesis, and termination occurs when specific conditions are satisfied; therefore its output must be correct. Theorem 3 indicates that REMAP learns the correct minimal automaton isomorphic to the target automaton in polynomial time.
Next, we show how Remap achieves PAC–identification when sampling-based equivalence queries are used.
Definition 5 (Probably Approximately Correct Identification).
Given Moore machine $M=\langle Q,q_{0},\Sigma^{I},\Sigma^{O},\delta,L\rangle$, let the classification function $f:(\Sigma^{I})^{*}\rightarrow\Sigma^{O}$ be represented by $f(s)=L(\delta(q_{0},s))$ for all $s\in(\Sigma^{I})^{*}$. Let $\mathcal{D}$ be any probability distribution over $(\Sigma^{I})^{*}$. An algorithm $\mathcal{A}$ probably approximately correctly identifies $f$ if and only if for any choice of $0<\epsilon\leq 1$ and $10<d<$ 1, $\mathcal{A}$ always terminates and outputs an $\epsilon$-approximate sequence classifier $\hat{f}:(\Sigma^{I})^{*}\rightarrow\Sigma^{O}$, such that with probability at least $1-d$, the probability of misclassification is $P(\hat{f}(s)\neq f(s))\leq\epsilon$ when $s$ is drawn according to the distribution $\mathcal{D}$.
Theorem 4.
Remap achieves probably approximately correct identification of any Moore machine when the teacher $\mathcal{T}$ uses sampling-based equivalence queries with at least $m_{k}\geq\left\lceil\frac{1}{\epsilon}\left(\ln\frac{1}{d}+k\ln 2\right)\right\rceil$ samples drawn i.i.d. from $\mathcal{D}$ for the $k$th equivalence query.
Theorem 5.
To achieve PAC-identification under Remap, given parameters $\epsilon$ and $d$, and if $f$ can be represented by a minimal Moore machine with $n$ states and $|\Sigma^{O}|$ classes, then teacher $\mathcal{T}$ needs to sample at least $\mathcal{O}\left(n+|\Sigma^{O}|+\frac{1}{\epsilon}\left((n+|\Sigma^{O}|)\ln\frac{1}{d}+(n+|\Sigma^{O}|)^{2}\right)\right)$ sequences i.i.d. from $\mathcal{D}$ over the entire run of Remap.
Theorem 4 and Theorem 5 imply Remap achieves PAC–identification for a choice of $\epsilon$ and $d$ as long as the teacher samples sufficient sequences per equivalence query.
Finally, Moore and Mealy machine equivalence permits converting the final hypothesis to a reward machine.
Learning Reward Machines from Preferences
We consider applying Remap to learn reward machines from preferences. Reward machines are Mealy machines with propositional and reward semantics. Equivalence between Mealy and Moore machines allows the output of Remap to be converted to a reward machine. We first review reinforcement learning and reward machine semantics.
Markov Decision Processes  Decision making problems are often modeled by a Markov Decision Process (MDP), which is a tuple $\langle\mathcal{S},\mathcal{A},P,R,\gamma\rangle$ where $\mathcal{S}$ is the set of states, $\mathcal{A}$ is the set of actions, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ represents the transition probability from state $s$ to $s^{\prime}$ via action $a$. The reward function $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ provides the associated scalar reward, and $0\leq\gamma\leq 1$ is a discount factor. In MDPs, the Markovian assumption is that transitions and rewards depend only upon the current state-action pair and the next state. However, Abel et al. (2021) show not all tasks are expressible using Markovian reward.
Non-Markovian Reward  Non-Markovian Reward Decision Processes (NMRDP) are identical to MDPs, except that $R:(\mathcal{S}\times\mathcal{A})^{*}\rightarrow\mathbb{R}$ is non-Markovian a reward function that depends on state-action history. This allows reward machines to model a class of non-Markovian reward functions.
Reward Machines  A reward machine (RM) is a Mealy machine where $\Sigma^{I}=2^{\mathcal{P}}$, $\Sigma^{O}$ is a set of reward emitting objects, and $\mathcal{P}$ is a set of propositions describing states and actions. A labeling function $\mathbb{L}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow 2^{\mathcal{P}}$ with $\mathbb{L}(s_{k-1},a_{k},s_{k})=l_{k}$ labels a state-action sequence $s_{0}a_{1}s_{1}a_{2}s_{2}\dots a_{n}s_{n}$ with label sequence $l_{1}l_{2}\dots l_{n}$. Thus, reward machines operate over label sequences.
A single disjunctive normal formula (DNF) labeled transition can be used to summarize multiple transitions with identical $\Sigma^{O}$ labels (connecting a pair of states), since the elements of $2^{\mathcal{P}}$ are sets of propositions. Reward machines map label sequences to reward outputs and can be represented as $f:(2^{\mathcal{P}})^{*}\rightarrow\Sigma^{O}$, so Remap learns a RM by converting the output Moore machine to a reward machine.
Sequential Tasks in OfficeWorld and CraftWorld  Icarte et al. (2018) and Andreas et al. (2017) introduced the OfficeWorld and CraftWorld gridworld domains, respectively, and feature sequential tasks encoded as reward machines. OfficeWorld features 4 sequential tasks across several rooms with various objects available for an agent to interact with. Example tasks include (1) picking up coffee and mail and delivering them to a certain room, or (2) continuously patrolling between a set of rooms. CraftWorld is a 2D version of MineCraft, where the 10 sequential tasks involve the agent collecting materials and constructing tools or objects in a certain order while avoiding hazards.
Empirical Results
We evaluate the exact and PAC–identification (PAC-ID) versions of Remap and consider: first, how often is PAC-ID Remap correct? (Exact Remap is guaranteed to be correct). To answer this, we run experiments by applying PAC-ID Remap to learn reward machines (RMs), by converting the Moore machine into a RM. We measure empirical correctness with empirical probability of isomorphism and average regret. Second, how do exact and PAC-ID Remap scale? We measure preference query complexity as a function of the number of unique sequences queried, and present an example phase diagram of algorithm execution. We use Z3 (De Moura and Bjørner 2008) for the constraint solver.
Setup  We investigate these questions on 14 sequential tasks in the OfficeWorld and CraftWorld domains. The Appendix A contains domain specific details. We implement exact equivalence queries using a variant of the Hopcroft-Karp algorithm (Almeida, Moreira, and Reis 2009; Hopcroft and Karp 1971). We implement i.i.d. sequence sampling in sampling-based equivalence queries with the following process per sample: sample a length $L$ from a geometric distribution; then, construct an $L$-length sequence by drawing $L$ elements i.i.d. from a uniform distribution over $\Sigma^{I}$.
PAC–Identification Correctness Experiments
Reproducibility  PAC-ID Remap was run 100 times per ground truth reward machine. We measure correctness based on (1) empirical probability that the learned RM is isomorphic to the ground truth RM, based on classification accuracy, and (2) average policy regret between the learned RM policy and the ground truth RM policy.
Figure 3: PAC–identification Remap: (left) empirical isomorphism probability, (right) average regret as functions of the number of samples per equivalence query for 4 OfficeWorld tasks (O-T1-4) and 10 CraftWorld tasks (C-T1-10).
Figure 4: Query Complexity. Top left: Exact Remap preference query complexity. Mean of 100 trials per ground truth reward machine (blue dots) ±$\pm 1$ standard deviation (orange, grey bars). Top right: PAC-ID Remap, preference query complexity is $\mathcal{O}(n\ln n)$ in the number of unique sequences in the table. Bottom left: Example termination phase diagram. Bottom right: CraftWorld environment depiction.
Empirical Probability of Isomorphism is the fraction of learned RMs with 100% classification accuracy. As the number of sample sequences tested by the teacher per equivalence query increases, the probability that the learner outputs a RM isomorphic to the ground truth RM upon termination goes to $1$ (Figure 3, left column). Classification accuracy is defined as the fraction of a test set of sequences that are identically classified by the learned and ground truth RMs. The Appendix describes the distribution over $(\Sigma^{I})^{*}$ that the test set is drawn from.
Average Policy Regret  We employ Q-learning with counterfactual experiences for reward machines (CRM) (Toro Icarte et al. 2022) to obtain optimal policies for ground truth and learned RMs. We measured the empirical expected return of optimal policies learned from each type of RM. Average regret for a given task was measured as the difference between the empirical return under the ground truth RM for that task (averaged over 100 CRM trials) and the empirical return under the learned RM (with 10 CRM trials per learned RM, then averaging all $100\times 10=1000$ trials). Regret goes to 0 as the number of samples tested by the teacher per equivalence query increases (Figure 3, right column). The Appendix A describes regret computation details.
Correctness Conclusion  Exact Remap learns the correct automaton 100% of the time. Additionally, PAC-ID Remap is more likely to be correct as the number of samples per equivalence query increases: isomorphism probability goes to 1 and regret goes to 0 for all tasks in both domains.
Scaling Experiments
Figure 4 shows query complexity results. We measure preference query complexity of exact and PAC-ID Remap, as a function of the number of unique sequences stored in the table upon termination. Exact Remap (upper left) displays a trendline of $C=0.2121N\ln N$ with $R^{2}=0.9963$, where $C$ is the number of queries, and $N$ is the number of unique sequences in the observation table. PAC-ID Remap (upper right) tends to make significantly more preference queries about unique sequences compared to exact Remap due to the sampling process. However, the number of preference queries is still $\mathcal{O}(N\ln N)$ due to randomized quicksort and linear merge comparison complexity.
The termination phase diagram in Figure 4 illustrates the taxi distance from $(0,1)$ to $(|\Sigma^{O}|,n)$ is the maximum number of equivalence queries Remap makes (Theorem 3). Progress (Lemma 1 and Theorem 2) towards termination (Corollary 1) occurs whenever a new hypothesis is made. Remap can terminate early when all variables have correct values and the required number of states is reached.
Related Work
Active approaches for learning automata are variations or improvements of Angluin’s seminal $\textrm{L}^{*}$ algorithm (1987), featuring membership and equivalence queries. We consider an alternative formulation: actively learning automata from preference and equivalence queries featuring feedback. We first discuss adaptations of $\textrm{L}^{*}$ for learning variants of finite automata, including reward machine specific variants.
Learning Finite Automata  Angluin (1987) introduced $\textrm{L}^{*}$ to learn deterministic finite automata (DFAs). Remap has similar theoretical guarantees as $\textrm{L}^{*}$, but utilizes a symbolic observation table, rather than an evidence-based one. Other algorithms adopt the evidence-based table of $\textrm{L}^{*}$ to learn: symbolic automata (Drews and D’antoni 2017; Argyros and D’antoni 2018), where Boolean predicates summarize state transitions; weighted automata (Bergadano and Varricchio 1994; Balle and Mohri 2015) which feature valuation semantics for sequences on non-deterministic automata; probabilistic DFAs (Weiss, Goldberg, and Yahav 2019), a weighted automata that models distributions of sequences. None of these approaches uses preference queries.
However, Shah et al. (2023) considers active, cost-based selection between membership and preference queries to learn DFAs, relying on a satisfiability encoding of the problem. They assume a fixed hypothesis space and have probabilistic guarantees for termination and correctness. Remap, through unification, navigates a sequence of hypothesis spaces, each guaranteed to contain a concrete hypothesis satisfying current constraints, and has theoretical guarantees of correctness, minimalism, and termination under exact and PAC–identification settings.
Furthermore, learning finite automata from preference information relates to the novel problem of learning reward machines (Icarte et al. 2018) from preferences. Learning Markovian reward functions from preferences has be studied extensively using neural (Bıyık, Talati, and Sadigh 2022; Sadigh et al. 2017) and interpretable decision tree (Bewley and Lécué 2021; Kalra and Brown 2023) representations, but approaches for learning reward machines primarily adapt evidence-based finite automata learning approaches.
Reward Machine Variants  Several reward machine (RM) variants have been proposed. Classical RMs (Icarte et al. 2018; Toro Icarte et al. 2019) have deterministic transitions and rewards; probabilistic RMs (Dohmen et al. 2022) model probabilistic transitions and deterministic rewards; and stochastic RMs (Corazza, Gavran, and Neider 2022) pair deterministic transitions with stochastic rewards. Symbolic RMs (Zhou and Li 2022) are deterministic like classical RMs, but feature symbolic reward values in place of concrete values. Zhou and Li (2022) apply Bayesian inverse reinforcement learning (BIRL) to infer optimal reward values and actualize symbolic RMs into classical RMs, and require a symbolic RM sketch. Remap requires no sketch, since it navigates over a hypothesis space of symbolic RMs and outputs a concrete classical RM upon termination.
Learning Reward Machines  Many RM learning algorithms assume access to explicit reward samples via environment interaction. Given a maximum RM size, Toro Icarte et al. (2019) apply discrete optimization to arrive at a perfect classical RM. Xu et al. (2020) learn a minimal classical RM by combining regular positive negative inference (Dupont 1994) with Q-learning for RMs (Icarte et al. 2018), and apply constraint solving to ensure each hypothesis RM is consistent with observed reward samples. Corazza et al. (2022) extended the method to learn stochastic RMs. Topper et al. (2024) extends BIRL to learn classical RMs using simulated annealing, but needs the number of states to be supplied, and requires empirical tuning of hyperparameters. $\textrm{L}^{*}$ based approaches have also been used to learn classical (Tappler et al. 2019; Gaon and Brafman 2020; Xu et al. 2021) and probabilistic (Dohmen et al. 2022) RMs, relying on concrete observation tables. Gaon and Brafman and Xu et al. use a binary observation table, while Tappler et al. and Dohmen et al. record empirical reward distribution table entries.
In contrast, Remap uses a symbolic observation table, and uses preferences information in place of explicit reward values. Remap navigates symbolic hypothesis space, with constraint solving enabling a concrete classical RM.
Conclusion
We introduce the problem of learning Moore machines from preferences and propose Remap, an $\textrm{L}^{*}$ based algorithm, wherein a strong learner with access to a constraint solver is paired with a weak teacher capable of answering preference queries and providing counterexample feedback in the form of a constraint. Unification applied to a symbolic observation table permits symbolic hypothesis space navigation; the constraint solver enables concrete hypotheses. Remap has theoretical guarantees for correctness, termination, and minimalism under both exact and PAC–identification settings, and it has been empirically verified under both settings when applied to learning reward machines. Next steps include investigating more realistic preference models, handling variable strength feedback, and handling inconsistency.
Appendix A Technical Appendix
In this technical appendix, we present empirical data, examples, experimental setup information, termination plots, detailed algorithms, and proofs. We also present the reproducibility checklist at the end of the appendix.
Experimental Details
Hardware
Experiments were run on a server with 512GB of memory and 4 Intel(R) Xeon(R) Gold 6248R CPUs. Each experimental run was executed on a single core. The operating system was Ubuntu 22.04.4. The relevant software libraries are listed in the code appendix remap/README file.
OfficeWorld and CraftWorld Domains
The computational experiments in this paper involved the OfficeWorld and CraftWorld domains. Both are gridworld environments that have been used in the reward machine literature. They feature sequential tasks which can be represented as a reward machine. Figure 5 illustrates an example OfficeWorld domain, along with a corresponding reward machine representing a task where the agent must bring paper to the desk while avoiding obstacles.
Handling Reward Machine Incompleteness.
The classical reward machines specified by Icarte et al. (2018) in the OfficeWorld and CraftWorld domains are incomplete automata, in that not all states have transitions defined. Specifically, classical reward machines have a terminal state from which no transitions can occur. However, an incomplete reward machine can be converted into a complete one by adding transitions from all terminal states to a single, special absorbing “HALT” state.
Therefore, to handle reward machine incompleteness, we first convert the incomplete reward machines into complete Mealy machines with terminal states, and then convert the Mealy machine to a complete Moore machine with a single absorbing HALT state which all terminal states transition to. This latter Moore machine is used by the teacher as the ground truth non-Markovian reward function. Next, Remap is run, resulting in a learned Moore machine. The learned Moore machine is converted to a Mealy machine, and the absorbing state is identified and removed.
We do not merge terminal states in order to remain consistent with the original implementation reward machines, but we do collapse pairs of transitions into single summary transitions if a pair of transitions share the same start, end, and output values. This is accomplished by constructing a truth table, constructing a disjunctive normal formula (DNF) for the truth table, and then simplifying the DNF and using the result as the summarizing transition label. Once this process is complete, the result is a learned reward machine, which can now be evaluated.
Handling Reward Machine Nondeterminism.
Classical reward machines have a deterministic definition. However, some of the classical reward machines specified by Icarte et al. (2018) are nondeterministic, in that from a given state, multiple transitions can be satisfied using under a given set of true propositions. This was the case for the reward machines for tasks 5 through 10 in the CraftWorld domain.
The common nondeterminism in those reward machines was the following type: assume the proposition set is $\mathcal{P}=\{a,b\}$, and the transitions have been summarized according to some set of Boolean formulae for which a subset is $\{\phi_{a},\phi_{b}\}$, where $\phi_{a}$ is satisfied whenever $a$ holds, and $\phi_{b}$ is satisfied whenever $b$ holds. Suppose we have a state $q_{1}$, and two of its summarized transitions are the following: $q_{2}=\delta(q_{1},\phi_{a})$ and $q_{3}=\delta(q_{1},\phi_{b})$, meaning that a transition from $q_{1}$ to $q_{2}$ occurs if $\phi_{a}$ is satisfied, and a transition from $q_{1}$ to $q_{3}$ occurs if proposition $\phi_{b}$ holds. Clearly, $\phi_{a}$ and $\phi_{b}$ can simultaneously be satisfied if $a\land b$ holds, implying nondeterminism. Additionally, all the nondeterministic reward machine specifications also contained the following style of transitions: $q_{4}=\delta(q_{1},\phi_{a}\phi_{b})=\delta(q_{1},\phi_{b}\phi_{a})$, where $\phi_{a}\phi_{b}$ and $\phi_{b}\phi_{a}$ are sequences of length 2. This type of nondeterminism can be corrected by adding an additional state $q_{5}$ and modifying the transitions from $q_{1}$ to $q_{2}$ and $q_{3}$, while still maintaining the intended behabior of reaching state $q_{4}$. Specifically, the conversion is, given
Figure 5: Example OfficeWorld domain (top) and a reward machine (bottom) encoding the sequential task “bring the paper to the desk without running into any obstacles.” A transition from state $u_{i}$ to $u_{j}$, labeled by propositional formula $\phi$ and scalar reward $r$ as $\langle\phi,r\rangle$, occurs only if $\phi$ is satisfied; reward $r$ is emitted upon transition. Propositions $p,t$, and $d$ are the agent: possessing paper, running into an obstacle, and being located at the desk.
$q_{1}$ $=\delta(q_{1},\varepsilon)=\delta(q_{1},\neg\phi_{a}\land\neg\phi_{b})$
$q_{2}$ $=\delta(q_{1},\phi_{a})=\delta(q_{2},\neg\phi_{b})$
$q_{3}$ $=\delta(q_{1},\phi_{b})=\delta(q_{3},\neg\phi_{a})$
$q_{4}$ $=\delta(q_{1},\phi_{a}\phi_{b})=\delta(q_{1},\phi_{b}\phi_{a}),$
we can make the following adjustments and additions:
$q_{2}=\delta(q_{1},\phi_{a})$ $\Longrightarrow q_{2}=\delta(q_{1},\phi_{a}\land\neg\phi_{b})$
$q_{3}=\delta(q_{1},\phi_{b})$ $\Longrightarrow q_{3}=\delta(q_{1},\phi_{b}\land\neg\phi_{a})$
Add new state $\Longrightarrow q_{5}=\delta(q_{1},\phi_{a}\land\phi_{b})$
[…]
Add new transition $\Longrightarrow q_{4}=\delta(q_{5},\phi_{a}\lor\phi_{b}).$
These changes make the reward machine deterministic while still maintaining desired behavior by explicitly providing three different transitions away from state $q_{1}$ for processing the input proposition sets $\{a\},\{b\},$ and $\{a,b\}$ separately.
Figure 6: Learning a Moore machine with Remap (top) vs $\textrm{L}^{*}$ (bottom). $\textrm{L}^{*}$ employs concrete values in its observation table, whereas Remap uses a symbolic observation table $\langle S,E,T;\mathcal{C},\Gamma\rangle$. The function $f(s)$ returns $1$ if $s$ is in $a^{*}b$ and returns 00 otherwise. (a) initializing $\langle S,E,T;\mathcal{C},\Gamma\rangle$ and performing a SymbolicFill, unclosed table; (b) expand $S$ with sequence $b$ to close the table, followed by a SymbolicFill yielding a unified, closed, and consistent $\langle S,E,T;\mathcal{C},\Gamma\rangle$; (c) MakeHypothesis yields concrete hypothesis $h_{1}$ from the symbolic hypothesis and solving constraints in $\mathcal{C}$; (d) submit $h_{1}$ via equivalence query, receive and process the counterexample $bab$ with feedback $f(bab)=0$ from teacher by adding $bab$, $ba$, and $b$ to $S$, perform a SymbolicFill, set the value equivalence class of $T(bab)$ to $f(bab)$, yields an inconsistent table; (e) expand suffixes $E$ with $b$ and perform a SymbolicFill, yielding a unified, closed, and consistent table; (f) MakeHypothesis yields concrete hypothesis $h_{2}$ for an equivalence query, wherein the teacher establishes $h_{2}$ is correct, and the algorithm terminates. Learning the target Moore machine with $\textrm{L}^{*}$ is shown is parts (g)-(l). (g) shows using membership queries to populate the table with concrete values, resulting in an unclosed table, which is then made closed in (h) by adding $b$ to $S$. Since the result is closed and consistent, a concrete hypothesis can be made in part (i). Sending this hypothesis to the teacher via an equivalence query results in the teacher sending a counterexample $bab$ back, which must be processed in (j). Here, the $bab$ and all its prefixes are added to $S$, resulting in an inconsistent table. Consistency is resolved by adding $b$ to $E$, resulting in a closed and consistent table. This allows the final hypothesis to be made in part $(l)$.
Test Set Generation for Classification Accuracy
Here, we describe the process inducing the distribution $\mathcal{D}$ over $(\Sigma^{I})^{*}$ from which the test set is generated. To generate the test set, we follow the procedure described for sampling-based equivalence queries: we first sample a random variable $L$ from a geometric distribution to represent desired sequence length, and then populate each of the $L$ sequence elements i.i.d. from the uniform distribution over $\Sigma^{I}$. The geometric distribution we use is $\text{Pr}(L=l)=(1-p)^{l-1}p$, where $p=0.2$. Thus, the average sampled sequence length is 5.
Next, we amend this sample sequence set with the set of sequences guaranteed to be composed only from explicitly specified transitions in the incomplete ground truth reward machine. This latter sequence sample set is generated by finding all sequences of Boolean formula from the initial state to all other states via iterative deepening search, resulting in all sequences with length at most the number of states in the reward machine. Each Boolean formula sequence generates multiple sequences with elements from $2^{\mathcal{P}}$, by uniformly sampling elements from $2^{\mathcal{P}}$ that make the formula true. Explicitly, for a given path of length $d$ from the start state to depth $d$, where $d$ ranges from $2$ to $N$ the number of states in the automaton, we generate $d|2^{\mathcal{P}}|s$ sequences, where $s$ is a positive integer. Thus, for an given automaton with $N$ states, we have at most $\mathcal{O}(|2^{\mathcal{P}}|^{N})$ paths, so we generate a test set of size $\mathcal{O}(|2^{\mathcal{P}}|^{N+1}ds)$. Each of the elements from the set $\{25,50,100,200,300,400,500,1000,2000,5000\}$ was used as the value of $s$.
For tasks 1 through 4 of OfficeWorld and CraftWorld, this corresponded to between 35.4k to 36.2k samples. For tasks 5 through 10 of CraftWorld, this corresponded to 1.4m to 24.6m samples.
We evaluate each ground truth and learned reward machine pair with these sequences. Classification accuracy is the fraction of sequences in the sample set identically classified by the learned and ground truth reward machines.
Regret Computation Details
For the regret experiments, we needed to train and evaluate policies under the ground truth and learned reward machines. For a given reward machine, REMAP was run 100 times, producing a set of 100 reward machines per ground truth. Any differences in reward machines stemmed from how the teacher sampled test sequences and the order and length of counterexamples presented.
To compute regret, we computed empirical regret between the optimal policy under the ground truth reward machine and the optimal policy of each of the learned reward machines.
We utilized Q-learning with counterfactual experiences for reward machines (CRM) (Toro Icarte et al. 2022) to obtain optimal policies for each reward machine. We set the discount factor to $\gamma=0.9$, and for the OfficeWorld reward machines, each policy was trained for a total of $2\times 10^{5}$ steps, and $2\times 10^{6}$ steps for CraftWorld reward machines. Each reward machine was trained for at least 10 seeds. Observing the return curves, we concluded that by $1\times 10^{5}$ steps, the policy was already optimal for OfficeWorld domains, and by $1\times 10^{6}$ steps for CraftWorld domains. We compute the average reward per step of the optimal policy by summing the total return of the policy over the last $10^{5}$ (OfficeWorld) or $10^{6}$ (CraftWorld) steps:
[…]
$=\hat{R}_{k}$
Empirical Average Reward per Step $=\frac{1}{N}\sum_{k=1}^{N}\hat{R}_{k}$
$=\hat{\mathbb{E}}[\hat{R}],$
where $r_{k,t}$ represents the reward received for seed $k$ at step $t$, and $s$ and $s+\Delta s$ represent the interval of steps the average is taken over, and $N$ represents the total number of seeds. Then, the average regret plotted in the paper was the difference between the empirical average reward per step of the optimal policy induced by the ground truth reward machine, and the empirical average reward per step of the optimal policies from the corresponding learned reward machines. Sample variance was computed for the ground truth reward machines via
$\text{Var}(\hat{R})=\frac{1}{N-1}\sum_{k=1}^{N}(\hat{R}_{k}-\hat{\mathbb{E}}[\hat{R}])^{2}$
with standard deviation computed via taking the square root of the sample variance.
Example Comparing REMAP and $\textrm{L}^{*}$
Figure 2 of the main paper contained an abridged example comparing REMAP and $\textrm{L}^{*}$. We show the full example in Figure 6 of the Appendix showing how REMAP and $\textrm{L}^{*}$ learn the Moore machine shown in part (l).
If the ground truth is to be interpreted as the Moore equivalent of a reward machine, then the ground truth Moore machine has 3 states, with $q_{b}$ as the terminal state, $q_{ba}$ is the absorbing HALT state, and $q_{\varepsilon}$ as the initial state; the corresponding ground truth reward machine would have only two states—$q_{\varepsilon}$ as the initial state, and $q_{b}$ as the terminal state, with no transitions out of $q_{b}$; there would be no absorbing $q_{ba}$ state. Additionally, the transition from $q_{\varepsilon}$ to $q_{b}$ via $b$ would have a reward of $1$ associated with it, and the self-transition $q_{\varepsilon}$ to $q_{\varepsilon}$ via $a$ would have a reward of 00 associated with it; the states would no longer have rewards associated with them.
If the input alphabet is interpreted as the powerset of a proposition set $\mathcal{P}=\{p\}$, with $2^{\mathcal{P}}=\Sigma^{I}=\{\{\},\{p\}\}$, then we can use $a=\{\}$ and $b=\{p\}$ for convenience, and where $\Sigma^{O}=\{0,1\}$. As shown in the example, REMAP uses a symbolic observation table and performs preference queries, and closedness and consistency tests are evaluated with respect to a unified table. $\textrm{L}^{*}$ uses a concrete observation table and uses membership queries.
Figure 7: Empirical Scaling Plots. Left to right: Under an inexact PAC-identification teacher, empirical distributions of how the total number of sequences in the table upon termination of Remap depends on alphabet size, number of states, and maximum counterexample length (color coded using data from CraftWorld tasks C-T5 through C-T10, see Figure 3 legend).
Additional Scaling Measurements
Since we measured query complexity as a function of the number of unique sequences present in the table upon termination, we also measured the number of unique sequences as functions of input alphabet size, number of states, and length of the maximum counterexample received. These results are shown in Figure 7.
Termination Plots
We present a full set of termination phase diagrams for learning the CraftWorld reward machines for tasks 5 through 10 in Figure 8. Columns correspond to tasks, while rows correspond to number of samples the teacher make per equivalence query. Each plot contains 100 paths through the termination phase space, where the $x$-axis is the number of known representative values, and the $y$-axis is the number of states in the hypothesis. Each node along the path corresponds with an event in Remap. The start of each path always starts at $(0,1)$, since there is always an initial state, but no known representatives. Green $X$’s represent when an equivalence query is made. Blue circles represent the immediate result of a closure operation, while orange squares represent the immediate result of a consistency operation. The red star represents the upper bound on number of states and number of known representative values. Observe that between consecutive equivalence queries, at least one of the number of states or the number of known representative values must increase: in particular, if the number of known representative values does not increase as the result of an equivalence query, the number of states must increase. Additionally, it is possible for Remap to terminate early, when the number of states has reached the upper bound, and when the satisfying solution to the constraints has all correct values.
Figure 8: All phase diagram termination plots for CraftWorld tasks 5 through 10. From left to right: Task 5, 6, 7, 8, 9, 10. From top to bottom: Teacher tests 25, 50, 100, 200, 300, 400, 500 samples per equivalence query. The X-axis is number of known representative values in the hypothesis, the Y-axis is the number of states in the hypothesis. Legend: blue circle is a closure operation, orange square is a consistency operation, green x is an equivalence query, and the red star represents the the upper bound termination condition. We observe it is possible to terminate early, prior to reaching the upper bound. Each plot contains 100 individual paths through phase space.
Remap Algorithms
In this section, we present the algorithms used in Remap. Algorithm 1 presented in the body of the paper is an abbreviated version of Algorithm 3 (Remap). In Algorithm 3, we expand the function MakeClosedAndConsistent into loop for making the symbolic observation table unified, closed, and consistent. We also present the Symbolic Fill Procedure in Algorithm 5, which is responsible for (1) creating fresh variables for sequences which have not been queried, (2) performing preference queries, and (3) performing unification. Algorithm 6 illustrates how preference queries are performed as comparisons, and also shows how the constraint set is updated with the return information. Algorithm 7 is responsible for constructing a hypothesis from a unified, closed, and consistent observation table, and it includes the FindSatisfyingSolution (abbreviated as FindSolution in the main paper) procedure which encodes all the collected constraints, known representative values, and global constraints, and sends them to the solver. Algorithm 8 illustrates the generic equivalence query used by the teacher for probably approximately correct (PAC) identification. Note that in practice, for both the equivalence query and for obtaining classification accuracy, we collect all sampled sequences into a set first, then perform evaluation over the set. For the equivalence query, the first counterexample encountered is returned (and the remainder is untested). For classification accuracy, we evaluate all the sample sequences. Finally, in Algorithm 9, we present the unification algorithm used in Remap. Overall, the algorithm has three sections: (1) creation and merging of equivalence classes and electing representatives, which converts all collected equality relations into equivalence classes (sets of variables which are equivalent), (2) performing unification on the collection of inequalities, performed via substituting each variable with its elected representative, and (3) replacing each variable in the symbolic observation table with the elected representative of the equivalence class of the variable.
Algorithm 3 Remap
Input: Input alphabet $\Sigma^{I}$, output alphabet $\Sigma^{O}$, and a teacher $\mathcal{T}$
Output: Moore Machine $\mathcal{H}=\langle\hat{Q},\Sigma^{I},\Sigma^{O},\hat{q}_{0},\hat{\delta},\hat{L}\rangle$
1:  Initialize observation table $(\langle S,E,T\rangle)$, constraint set $\mathcal{C}$, and context $\Gamma$ as $\langle S,E,T;\mathcal{C},\Gamma\rangle$ with $S=\{\varepsilon\},E=\{\varepsilon\}$, $\mathcal{C}=\{\}$, $\Gamma=\emptyset$
2:  $\langle S,E,T;\mathcal{C},\Gamma\rangle\longleftarrow$ SymbolicFill$\left(\langle S,E,T;\mathcal{C},\Gamma\rangle|\mathcal{T}\right)$
3:  repeat
4:     while $\langle S,E,T;\mathcal{C},\Gamma\rangle$ is not closed and consistent do
5:        if $\langle S,E,T;\mathcal{C},\Gamma\rangle$ not consistent then
6:           Find $s_{1},s_{2}\in S$, $\sigma\in\Sigma^{I}$, and $e\in E$s.t. $\textbf{{row}}(s_{1})\equiv\textbf{{row}}(s_{2})$ and $\textbf{{row}}(s_{1}\cdot\sigma)\not\equiv\textbf{{row}}(s_{2}\cdot\sigma)$ and $T(s_{1}\cdot\sigma\cdot e)\not\equiv T(s_{2}\cdot\sigma\cdot e)$
7:           Add $\sigma\cdot e$ to $E$ via $E:=E\cup\{\sigma\cdot e\}$
8:        end if
9:        if $\langle S,E,T;\mathcal{C},\Gamma\rangle$ not closed then
10:           Find $s\in S$ and $\sigma\in\Sigma^{I}$ s.t. $\textbf{{row}}(s\cdot\sigma)\in\textbf{{rows}}(S\cdot\Sigma^{I})$ and $\textbf{{row}}(s\cdot\sigma)\notin\textbf{{rows}}(S)$
11:           Add $s\cdot\sigma$ to $S$ via $S:=S\cup\{s\cdot\sigma\}$
12:        end if
13:        $\langle S,E,T;\mathcal{C},\Gamma\rangle\longleftarrow$ SymbolicFill$\left(\langle S,E,T;\mathcal{C},\Gamma\rangle|\mathcal{T}\right)$
14:     end while
15:     $\mathcal{H}\longleftarrow$MakeHypothesis$\left(\langle S,E,T;\mathcal{C},\Gamma\rangle,\Sigma^{I},\Sigma^{O}\right)$
16:     result $\longleftarrow$ EquivalenceQuery$(\mathcal{H}|\mathcal{T})$
17:     if result $\neq$ correct then
18:        $(s^{\prime},r)\longleftarrow$ result
19:        for all $t\in\text{{{prefixes}}}(s^{\prime})$ do
20:           Add $t$ to $S$ via $S:=S\cup\{t\}$
21:        end for
22:        $\langle S,E,T;\mathcal{C},\Gamma\rangle\longleftarrow$ SymbolicFill$\left(\langle S,E,T;\mathcal{C},\Gamma\rangle|\mathcal{T}\right)$
23:        Update $\mathcal{C}$ to include constraint via $\mathcal{C}:=\mathcal{C}\cup\{\text{{GetVar}$(s^{\prime},\langle S,E,T;\mathcal{C},\Gamma\rangle)$}=r\}$
24:     end if
25:  until result $=$ correct
26:  return  $\mathcal{H}$
Algorithm 4 A Query Efficient Symbolic Fill Procedure
procedure SymbolicFill$\left(\langle S,E,T;\mathcal{C},\Gamma\rangle|\mathcal{T}\right)$
[…]
2:  for all $s\cdot e\in(S\cup(S\cdot\Sigma^{I}))\cdot E$ do
3:     if $s\cdot e\notin\Gamma$ then
4:        $\Gamma\left[s\cdot e\right],seqs\longleftarrow$ FreshVar(), $\text{{seqs}}\cup\{s\cdot e\}$
5:        $T(s\cdot e)\longleftarrow$ $\Gamma\left[s\cdot e\right]$
6:     end if
7:  end for
8:  sortseqs, $\mathcal{O}$longleftarrow PrefQsViaRandQuicksort(seqs, $\mathcal{O}|\mathcal{T}$)
[…]
10:  $\mathcal{O}\longleftarrow$ Unification$(\mathcal{O})$
11:  return  $\mathcal{O}$
Algorithm 5 A Query Inefficient Symbolic Fill Procedure
procedure SymbolicFill$\left(\langle S,E,T;\mathcal{C},\Gamma\rangle|\mathcal{T}\right)$
[…]
2:  for all $s\cdot e\in(S\cup(S\cdot\Sigma^{I}))\cdot E$ do
3:     if $s\cdot e\notin\Gamma$ then
4:        $\Gamma\left[s\cdot e\right]\longleftarrow$ FreshVar()
5:        $T(s\cdot e)\longleftarrow\Gamma\left[s\cdot e\right]$
6:        $\text{{newentries}}\longleftarrow\text{{newentries}}\cup\{s\cdot e\}$
7:     else
8:        $\text{{oldentries}}\longleftarrow\text{{oldentries}}\cup\{s\cdot e\}$
9:     end if
10:  end for
11:  for all $(s_{1}\cdot e_{1}),(s_{2}\cdot e_{2})\in$ PairCombinations(newentries, newentries $\cup$ oldentriesdo
12:     $p\longleftarrow$ PrefQuery$(s_{1}\cdot e_{1},s_{2}\cdot e_{2}|\mathcal{T})$
13:     $\mathcal{C}\longleftarrow$ UpdateConstraintSet$(p,s_{1}\cdot e_{1},s_{2}\cdot e_{2};\langle S,E,T;\mathcal{C},\Gamma\rangle)$
14:  end for
15:  $\langle S,E,T;\mathcal{C},\Gamma\rangle\longleftarrow$ Unification$(\langle S,E,T;\mathcal{C},\Gamma\rangle)$
16:  return  $\langle S,E,T;\mathcal{C},\Gamma\rangle$
Algorithm 6 Preference Query Procedure and Constraints Update Procedure
Preference queries are checked by the teacher executing each sequence using the ground truth transition function $\delta$ and checking the ground truth output value using $L$.
procedure PrefQuery$(s_{1},s_{2}|\mathcal{T})$
1:  if $L(\delta(q_{0},s_{1}))=L(\delta(q_{0},s_{2}))$ then
2:     return  00
3:  else if $L(\delta(q_{0},s_{1}))<L(\delta(q_{0},s_{2}))$ then
4:     return  $-1$
5:  else
6:     return  $+1$
7:  end if
procedure UpdateConstraintSet$(p,s_{1}\cdot e_{1},s_{2}\cdot e_{2};\langle S,E,T;\mathcal{C},\Gamma\rangle)$
1:  if $p=0$ then
2:     $\mathcal{C}:=\mathcal{C}\cup\left\{T(s_{1}\cdot e_{1})=T(s_{2}\cdot e_{2})\right\}$
3:  else if $p=-1$ then
4:     $\mathcal{C}:=\mathcal{C}\cup\left\{T(s_{1}\cdot e_{1})<T(s_{2}\cdot e_{2})\right\}$
5:  else
6:     $\mathcal{C}:=\mathcal{C}\cup\left\{T(s_{1}\cdot e_{1})>T(s_{2}\cdot e_{2})\right\}$
7:  end if
8:  return  $\mathcal{C}$
Algorithm 7 Make Hypothesis from Observation Table Procedure
procedure MakeHypothesis$\left(\langle S,E,T;\mathcal{C},\Gamma\rangle,\Sigma^{I},\Sigma^{O}\right)$
1:  $\hat{Q}=\{\textbf{{row}}(s)|\forall s\in S\}$ is the set of states
2:  $\hat{q}_{0}=\textbf{{row}}(\varepsilon)$ is the initial state
3:  $\hat{\delta}(\textbf{{row}}(s),\sigma)=\textbf{{row}}(s\cdot\sigma)$ for all $s\in S$ and $\sigma\in\Sigma^{I}$
4:  $\Lambda\longleftarrow$FindSatisfyingSolution$(\langle S,E,T;\mathcal{C},\Gamma\rangle,\Sigma^{O})$
5:  $\hat{L}(\textbf{{row}}(s))=\Lambda[T(s\cdot\varepsilon)]$ is the sequence to output function
6:  return  $\langle\hat{Q},\Sigma^{I},\Sigma^{O},\hat{q}_{0},\hat{\delta},\hat{L}\rangle$
procedure FindSatisfyingSolution$(\langle S,E,T;\mathcal{C},\Gamma\rangle,\Sigma^{O})$
1:  $\mathcal{C}_{val}:=$ Select$(\mathcal{C},Var\rightarrow Value)$
2:  $D=\left\{\left.\displaystyle\bigvee_{r\in\Sigma^{O}}T(s\cdot e)=r\right|\forall s\cdot e\in(S\cup(S\cdot\Sigma^{I}))\cdot E\text{ such that }\mathcal{C}_{val}[T(s\cdot e)]\equiv\bot\right\}$
3:  Submit to an SMT solver the constraint set $\mathcal{C}\cup D$
4:  return  The model $\Lambda$ which maps variables to values.
Algorithm 8 Sampling-based Equivalence Query
Equivalence is checked through sampling by the teacher. We assume a probability distribution $\mathcal{D}$ over all possible sequences $\left(\Sigma^{I}\right)^{*}$, and that the teacher can sample random sequences $s\sim\mathcal{D}$.
procedure EquivalenceQuery$\left(\langle\hat{Q},\Sigma^{I},\Sigma^{O},\hat{q}_{0},\hat{\delta},\hat{L}\rangle|\mathcal{T}\right)$
1:  repeat
2:     Sample a random sequence $s\sim\mathcal{D}$.
3:     if $L\left(\delta\left(q_{0},s\right)\right)\neq\hat{L}\left(\hat{\delta}\left(\hat{q_{0}},s\right)\right)$ then
4:        return  $\left(s,L\left(\delta\left(q_{0},s\right)\right)\right)$
5:     end if
6:  until up to $r$ times
7:  return  correct
Algorithm 9 Unification Procedure
Unification is performed by computing equivalence classes.
procedure Unification$\left(\langle S,E,T;\mathcal{C},\Gamma\rangle\right)$
1:  $\mathcal{C}_{ec}:=$ Select$(\mathcal{C},Var\rightarrow EquivalenceClass)$
2:  $\mathcal{C}_{val}:=$ Select$(\mathcal{C},Var\rightarrow Value)$
3:  $\mathcal{C}_{eq}:=$ Select$(\mathcal{C},Var=Var)$
4:  $\mathcal{C}_{in}:=$ Select$(\mathcal{C},Var<Var)\cup\text{{Select}}(\mathcal{C},Var>Var)$
5:  while $|\mathcal{C}_{eq}|>0$ do
6:     match Pop$(\mathcal{C}_{eq})$ with $L=R$
7:     if $L\in\mathcal{C}_{ec}\land R\in\mathcal{C}_{ec}$ then
8:        if $\mathcal{C}_{ec}\left[L\right]\not\equiv\mathcal{C}_{ec}[R]$ then
9:           $L_{rep}\longleftarrow$ GetRepresentative$(\mathcal{C}_{ec}\left[L\right])$
10:           $R_{rep}\longleftarrow$ GetRepresentative$(\mathcal{C}_{ec}\left[R\right])$
11:           Update $\mathcal{C}_{ec}\left[L\right]:=\mathcal{C}_{ec}\left[L\right]\cup\mathcal{C}_{ec}\left[R\right]$
12:           for all $v\in\mathcal{C}_{ec}\left[L\right]$ do
13:              Set $\mathcal{C}_{ec}\left[v\right]:=\mathcal{C}_{ec}\left[L\right]$
14:           end for
15:           if $L_{rep}\not\equiv R_{rep}\land R_{rep}\in\mathcal{C}_{val}$ then
16:              Remove redundant representative via Del$(\mathcal{C}_{val}\left[R_{rep}\right])$
17:           end if
18:        end if
19:     else if $L\in\mathcal{C}_{ec}\land R\not\in\mathcal{C}_{ec}$ then
20:        Update $\mathcal{C}_{ec}\left[L\right]:=\mathcal{C}_{ec}\left[L\right]\cup\left\{R\right\}$ and set $\mathcal{C}_{ec}\left[R\right]:=\mathcal{C}_{ec}\left[L\right]$
21:     else if $L\not\in\mathcal{C}_{ec}\land R\in\mathcal{C}_{ec}$ then
22:        Update $\mathcal{C}_{ec}\left[R\right]:=\mathcal{C}_{ec}\left[R\right]\cup\left\{L\right\}$ and set $\mathcal{C}_{ec}\left[L\right]:=\mathcal{C}_{ec}\left[R\right]$
23:     else
24:        Set $\mathcal{C}_{ec}\left[L\right]\longleftarrow$ EquivalenceClass($(\{L,R\})$ and $\mathcal{C}_{ec}\left[R\right]:=\mathcal{C}_{ec}\left[L\right]$
25:        Set $\mathcal{C}_{val}\left[\text{{GetRepresentative}}\left(\mathcal{C}_{ec}\left[L\right]\right)\right]\longleftarrow\bot$
26:     end if
27:  end while
[…]
29:  while $|\mathcal{C}_{in}|>0$ do
30:     constraint $\longleftarrow$ Pop$(\mathcal{C}_{in})$
[…]
32:        if $v\not\in\mathcal{C}_{ec}$ then
33:           Set $\mathcal{C}_{ec}\left[v\right]\longleftarrow$ EquivalenceClass($(\{v\})$
34:           Set $\mathcal{C}_{val}\left[\text{{GetRepresentative}}\left(\mathcal{C}_{ec}\left[v\right]\right)\right]\longleftarrow\bot$
35:        end if
[…]
37:     end for
[…]
39:  end while
40:  $\mathcal{C}_{in}:=$ subineqs
41:  for all $s\cdot e\in(S\cup(S\cdot\Sigma^{I}))\cdot E$ do
42:     Update $T[s\cdot e]$ via $T[s\cdot e]:=\text{{GetRepresentative}}\left(\mathcal{C}_{ec}\left[T[s\cdot e]\right]\right)$
43:  end for
44:  for all $s\cdot e\in\Gamma$ do
45:     Update $\Gamma[s\cdot e]$ via $\Gamma[s\cdot e]:=\text{{GetRepresentative}}\left(\mathcal{C}_{ec}\left[\Gamma[s\cdot e]\right]\right)$
46:  end for
Technical Proofs
We first establish the correctness of constructing a hypothesis deterministic finite automaton from a symbolic observation table via Algorithm 7, in Theorem 1 via lemmas 1-5. The proof strategies for lemmas 3 through 5 generally follows that of Angluin (1987), with appropriate adjustments to account for a set of possible hypotheses, compared to just a single hypothesis.
Lemma 1.
Let $\mathcal{H}$ be the set of all hypotheses that can be returned from
MakeHypothesis$(\langle S,E,T;\mathcal{C},\Gamma\rangle)$. All hypotheses in $\mathcal{H}$ are structurally identical. Two hypotheses $h_{1}$ and $h_{2}$ are structurally identical if they have equivalent states and equivalent transitions.
Proof.
We first show that all hypotheses in $\mathcal{H}$ must be structurally identical by construction, but have different $\hat{L}$ functions. Identical structure for all hypotheses in $\mathcal{H}$ can be shown by observing that lines 1-3 of Algorithm 7 are the same for each possible hypothesis in $\mathcal{H}$ for a given $\langle S,E,T;\mathcal{C},\Gamma\rangle$ input. Therefore, the only differences in the output hypothesis are due to lines 4 and 5. Since lines 1-3 construct the set of states and the transition function for each hypothesis, each hypothesis in $\mathcal{H}$ must therefore be structurally identical to all other hypotheses in $\mathcal{H}$. Lines 4 and 5 together select a satisfying solution for the free variables in $\langle S,E,T;\mathcal{C},\Gamma\rangle$, subject to the constraints $\mathcal{C}$. Therefore, each unique hypothesis in $\mathcal{H}$ is corresponds to a unique satisfying solution; the mapping between unique satisfying solutions and unique hypotheses is bijective. ∎
Lemma 2.
States in a Moore machine are represented by sets of sequences.
Proof.
Suppose we have a finite alphabet $\Sigma$, and a Moore machine, with set of states $Q$, initial state $q_{0}\in Q$, and transition function $\delta:Q\times\Sigma\rightarrow Q$. We define the initial state $q_{0}$ of the Moore machine to correspond with the empty sequence $\varepsilon$ which has length 00, so in fact we have $q_{0}=\delta(q_{0},\varepsilon)$, and for any $q\in Q$, $q=\delta(q,\varepsilon)$. The transition function can be extended recursively to $\delta:Q\times\left(\Sigma^{*}\right)\rightarrow Q$ by observing that $\delta(q_{0},\sigma\cdot\omega)=\delta(\delta(q_{0},\sigma),\omega)$, where $\sigma$ is a sequence of length of 1, and $\omega$ is a sequence of length of at least $1$. Similarly, $\delta(q_{0},\omega\cdot\sigma)=\delta(\delta(q_{0},\omega),\sigma)$. Based on this extended transition function, we can now consider the set $Q^{\prime}=\{\delta(q_{0},\sigma)|\sigma\in\Sigma\}$ to be a subset of $Q$.
First, we consider the case where $|Q^{\prime}|=|\Sigma|$, where all transitions from $q_{0}$ have led to unique states. We can therefore construct a mapping: $M:2^{\Sigma}\rightarrow Q^{\prime}$. We know that $|2^{\Sigma}|=2^{|\Sigma|}>|Q^{\prime}|$ if $|Q^{\prime}|=|\Sigma|\geq 1$, so we can consider a subset $K\subset 2^{\Sigma}$ such that $M$ is bijective. By constructing $K$, we can show that the elements of $K$ uniquely correspond to the elements of $Q^{\prime}$ because $M:K\rightarrow Q^{\prime}$ will be bijective. We note that if we construct $K=\{\{\sigma\}|\sigma\in\Sigma\}$, then clearly $|K|=|\Sigma|=|Q^{\prime}|$, satisfying the current case. Furthermore, the mapping $M(\{\sigma\})=\delta(q_{0},\sigma)\forall\sigma\in\Sigma$ is clearly a bijective mapping between $K$ and $Q^{\prime}$.
Next, we consider the case where $|Q^{\prime}|<|\Sigma|$, which implies that there exists at least one pair $\sigma_{i},\sigma_{j}\in\Sigma$ which lead to the same state (via the pidgeonhole principle). That is, $\delta(q_{0},\sigma_{i})=\delta(q_{0},\sigma_{j})$. If this is the case, then in order to make $M$ bijective, let $K=\{k|k=\bigcup_{i,j}\{\sigma_{i}\}\cup\{\sigma_{j}\}\forall(\sigma_{i},\sigma_{j})\in\Sigma\times\Sigma\text{ such that }\delta(q_{0},\sigma_{i})=\delta(q_{0},\sigma_{j})\}$. The following mapping for $M$ is bijective: $M(k)=\delta(q_{0},\sigma)\forall k\in K\text{ and }\forall\sigma\in k$, and that $K\subseteq 2^{\Sigma}$. Note each element of $K$ is a set of sequences of length 1.
Now, consider $Q^{\prime}_{n}=\{\delta(q_{0},\omega)|\omega\in(\Sigma)^{n}\}$ where $(\Sigma)^{n}$ denotes sequences of length at most $n$. Assume we can construct a bijective mapping $M:2^{(\Sigma)^{n}}\longrightarrow Q^{\prime}_{n}$ for all $1\leq n\leq N$ for some fixed $N$. We have already shown this for $N=1$ above. We now proceed to show that we can also construct a bijective mapping $M:2^{(\Sigma)^{N+1}}\longrightarrow Q^{\prime}_{N+1}$. Similar to the $N=1$ case, consider the case for when all sequences $\omega\in(\Sigma)^{N+1}$ lead to unique states—that is, when $|Q^{\prime}_{N+1}|=|\Sigma|^{N+1}$. Then the bijective mapping in this case is $M:K\longrightarrow Q^{\prime}_{N+1}$ where $K=\{\{\omega\}|\omega\in(\Sigma)^{N+1}\}$, and where $M(\{\omega\})=\delta(q_{0},\omega)$ for all $\omega\in(\Sigma)^{N+1}$, and $K\subset 2^{(\Sigma)^{N+1}}$.
Now, if $|Q^{\prime}_{N+1}|<|\Sigma|^{N+1}$, then by the pidgeonhole principle, multiple $\omega$ must lead to the same state. Hence, let $K=\{k|k=\bigcup_{i,j}\{\omega_{i}\}\cup\{\omega_{j}\}\forall(\omega_{i},\omega_{j})\in(\Sigma)^{N+1}\times(\Sigma)^{N+1}\text{ such that }\delta(q_{0},\omega_{i})=\delta(q_{0},\omega_{j})\}$. Then the bijective function is $M:K\longrightarrow Q^{\prime}_{N+1}$, where $M(k)=\delta(q_{0},\omega)$ for all $k\in K$ and for all $\omega\in k$, and note that $K\subseteq 2^{(\Sigma)^{N+1}}$.
Since we can construct bijective mappings from $M:2^{(\Sigma)^{n}}\longrightarrow Q^{\prime}_{n}$ for all $1\leq n\leq N$, and we have also shown this is the case for $n=N+1$, this must now also hold for all $n=N+d$, for integer $d\geq 0$, and thus this holds as $n$ tends towards infinity. Thus, states in a Moore machine are represented by sets of sequences, where each state corresponds to a set in $K$. Each state is therefore synonymous with an equivalence class of sequences—a set of sequences that are equivalent according to the extended transition function. ∎
Lemma 3.
The rows of a unified, closed, and consistent $\langle S,E,T;\mathcal{C},\Gamma\rangle$ represent the states in a Moore machine consistent with the constraints $\mathcal{C}$. The hypothesis MakeHypothesis$(\langle S,E,T;\mathcal{C},\Gamma\rangle)$ satisfies $\delta(q_{0},s)=\textbf{row}(s)$ for all $s\in(S\cup(S\cdot\Sigma))$.
Proof.
We utilize the inductive proof of Angluin (1987). For the case of $s=\varepsilon$, with length 00, we have $\delta(q_{0},\varepsilon)=q_{0}=\textbf{row}(\varepsilon)$, which is true by definition. Now, let us assume that $\delta(q_{0},s)=\textbf{row}(s)$ holds for all $s\in(S\cup(S\cdot\Sigma))$ with lengths no greater than $k$. Let $s^{\prime}\in(S\cup(S\cdot\Sigma))$ be a sequence of length $k+1$ such that $=s\cdot\sigma$. If $s^{\prime}\in S$, then $s\in S$ because $S$ is prefix closed. If $s^{\prime}\in S\cdot\Sigma$, then $s\in S$. We can then show that $\delta(q_{0},s^{\prime})=\delta(q_{0},s\cdot\sigma)=\delta(\delta(q_{0},s),\sigma)=\delta(\textbf{row}(s),\sigma)=\textbf{row}(s\cdot\sigma)=\textbf{row}(s^{\prime})$. ∎
Lemma 4.
The entries of a unified, closed, and consistent symbolic observation table correspond to sequence classification consistent with constraints $\mathcal{C}$. Specifically, for a unified, closed, and consistent $\langle S,E,T;\mathcal{C},\Gamma\rangle$, let $\Sigma^{*}$ be partitioned into at most $k\leq|\Sigma^{O}|$ disjoint sets $F_{1},F_{2},...,F_{k}$, and let $\Sigma^{O}_{k}$ be a specific subset of $k$ distinct elements of $\Sigma^{O}$. A sequence $s\in\Sigma^{*}$ is a member of $F_{j}$ if and only if it is classified as $\sigma_{j}\in\Sigma^{O}_{k}$. Then hypothesis MakeHypothesis$(\langle S,E,T;\mathcal{C},\Gamma\rangle)$ satisfies $\delta(q_{0},s\cdot e)\in F_{j}$ if and only if $T(s\cdot e)=\sigma_{j}$ for every $s\in(S\cup(S\cdot\Sigma))\text{ and every }e\in E$.
Proof.
We adapt the inductive proof of Angluin (1987), but make appropriate adjustments for sequence classification. Let $s\in(S\cup(S\cdot\Sigma))$ and $e=\varepsilon$ be the base case. Clearly, $\delta(q_{0},s\cdot\varepsilon)=\delta(q_{0},s)=\textbf{row}(s)$ as shown previously. If $s\in S$, then $\textbf{row}(s)\in F_{j}$ if and only if $T(s)=\sigma_{j}$. If $s\in S\cdot\Sigma$, then $\textbf{row}(s)\in\hat{Q}$, since the symbolic observation table is closed. A $\textbf{row}(s^{\prime})\in\hat{Q}$ is in $F_{j}$ if and only if $T(s^{\prime})=\sigma_{j}$.
Next, without loss of generality, assume that $\delta(q_{0},s\cdot e)\in F_{j}$ if and only if $T(s\cdot e)=\sigma_{j}$ for all sequences $e$ with length at most $k$. Let $e^{\prime}\in E$ be a sequence with length $k+1$ and let $s\in S\cup(S\cdot\Sigma)$. The sequence $e^{\prime}=\sigma\cdot e_{0}$ for some $e_{0}\in E$ of length $k$ and some $\sigma\in\Sigma$, because $E$ is suffix-closed. Furthermore, there is a sequence $s_{0}\in S$ such that $\textbf{row}(s)=\textbf{row}(s_{0})$ because the symbolic observation table is closed. Next, we show that if two sequences share a common suffix, but have different prefixes that end at the same state, then the two sequences also end at the same state. We observe that $\delta(q_{0},s\cdot e^{\prime})=\delta(q_{0},s\cdot\sigma\cdot e_{0})=\delta(\delta(q_{0},s),\sigma\cdot e_{0})=\delta(\textbf{row}(s),\sigma\cdot e_{0})=\delta(\textbf{row}(s_{0}),\sigma\cdot e_{0})=\delta(\delta(\textbf{row}(s_{0}),\sigma)e_{0})=\delta(\textbf{row}(s_{0}\cdot\sigma),e_{0})=\delta(\delta(q_{0},s_{0}\cdot\sigma),e_{0})=\delta(q_{0},s_{0}\cdot\sigma\cdot e_{0})$. This indicates that the sequences $s\cdot e^{\prime}$ and $s_{0}\cdot e^{\prime}$ share a common suffix, but potentially have different prefixes, the two sequences end at the same state because their prefixes end at the same state. Since $e_{0}$ has length $k$, and $s_{0}\cdot\sigma\in S\cup(S\cdot\Sigma)$, we can use our initial assumption that $\delta(q_{0},s_{0}\cdot\sigma\cdot e_{0})\in F_{j}$ if and only if $T(s_{0}\cdot\sigma\cdot e_{0})=\sigma_{j}$. Because $\textbf{row}(s)=\textbf{row}(s_{0})$, this means that $\textbf{row}(s)(e)=\textbf{row}(s_{0})(e)\forall e\in E$. Note that by definition, $T(s\cdot e^{\prime})=\textbf{row}(s)(e^{\prime})$, so $\textbf{row}(s_{0})(\sigma\cdot e_{0})=\textbf{row}(s)(\sigma\cdot e_{0})$ implies $T(s_{0}\cdot\sigma\cdot e_{0})=T(s\cdot\sigma\cdot e_{0})=T(s\cdot e^{\prime})$, which means $\delta(s\cdot e^{\prime})\in F_{j}$ if and only if $T(s\cdot e^{\prime})=\sigma_{j}$. ∎
Lemma 5.
Suppose $\langle S,E,T;\mathcal{C},\Gamma\rangle$ is a unified, closed, and consistent symbolic observation table. If the hypothesis $\hat{h}=\langle\hat{Q},\Sigma^{I},\Sigma^{O},\hat{q}_{0},\hat{\delta},\hat{L}\rangle$ from MakeHypothesis$(\langle S,E,T;\mathcal{C},\Gamma\rangle)$ has $n$ states, and $\hat{h}$ was generated using a satisfying solution $\Lambda$ to constraints $\mathcal{C}$, then any other Moore machine $h=\langle Q,\Sigma^{I},\Sigma^{O},q_{0},\delta,L\rangle$ with $n$ or fewer states that is also consistent with $\Lambda$ and $T$ is isomorphic to $\hat{h}$.
Proof.
We make additions and appropriate adjustments to the proof of Angluin (1987). First, let us partition $\hat{Q}$ into $k=|\Sigma^{O}|$ disjoint sets $\hat{F}_{1},...,\hat{F}_{k}$ such that for all $\sigma_{j}\in\Sigma^{O}$, $\hat{q}\in\hat{F}_{j}$ if and only if $\hat{L}(\hat{q})=\sigma_{j}$. Note that it is possible for some of the $\hat{F}_{j}$ to be empty. Similarly, let us also partition $Q$ into $k$ disjoint sets $F_{1},...,F_{k}$ such that for all $\sigma_{j}\in\Sigma^{O}$, $q\in F_{j}$ if and only if $L(q)=\sigma_{j}$. Next, we can define for all $q\in Q$ and for all $\sigma_{j}\in\Sigma^{O}$ the function $\textbf{row}^{\wedge}(q):E\rightarrow\Sigma^{O}$ such that $\textbf{row}^{\wedge}(q)(e)=\sigma_{j}$ if and only if $\delta(q,e)\in F_{j}$.
Because $h=\langle Q,\Sigma^{I},\Sigma^{O},q_{0},\delta,L\rangle$ is consistent with $\Lambda$ and $T$, we know that for each $s\in S\cup(S\cdot\Sigma)$ and for each $e\in E$, $\delta(q_{0},s\cdot e)\in F_{j}$ if and only if $T(s\cdot e)=\sigma_{j}$. Therefore, for all $e\in E$, $\textbf{row}^{\wedge}(q)(e)=\textbf{row}(s)(e)$ if $q=\delta(q_{0},s)$ implies $\textbf{row}^{\wedge}(\delta(q_{0},s))\equiv\textbf{row}(s)$. Next, because by definition $\hat{Q}=\{\textbf{row}(s)|\forall s\in S\}$, we have via substitution $\hat{Q}=\{\textbf{row}^{\wedge}(\delta(q_{0},s))|\forall s\in S\}$, which implies that $|Q|\geq n=|\hat{Q}|$. This is because in order for $\{\textbf{row}^{\wedge}(\delta(q_{0},s))|\forall s\in S\}$ to contain $n$ elements, $\delta(q_{0},s)$ must range over at least $n$ states as $s$ ranges over $S$. However, because we have assumed that $h$ contains $n$ or fewer states, $h$ must contain exactly $n=|\hat{Q}|=|Q|$ states.
Next, we can consider bijective mappings between $\hat{Q}$ and $Q$. Since $Q$ and $\hat{Q}$ have the same cardinality, and because $\textbf{row}^{\wedge}(\delta(q_{0},s))\equiv\textbf{row}(s)$, we know that for every $s\in S$ corresponds to a unique $q\in Q$, specifically, $q=\delta(q_{0},s)$. We can define the bijective mapping $\textbf{row}^{-\wedge}:\hat{Q}\rightarrow Q$ via for all $s\in S$, $\textbf{row}^{-\wedge}(\textbf{row}(s))=\delta(q_{0},s)$. From this mapping, we observe that $\textbf{row}^{-\wedge}(\textbf{row}(\varepsilon))=\textbf{row}^{-\wedge}(\hat{q}_{0})=\delta(q_{0},\varepsilon)=q_{0}$. Additionally, for all $s\in S$ and for all $\sigma\in\Sigma$, $\textbf{row}^{-\wedge}(\hat{\delta}(\textbf{row}(s),\sigma))=\textbf{row}^{-\wedge}(\textbf{row}(s\cdot\sigma))=\delta(q_{0},s\cdot\sigma)=\delta(\delta(q_{0},s),\sigma)=\delta(\textbf{row}^{-\wedge}(\textbf{row}(s)),\sigma)$, which implies that for all $s\in S$ and $\sigma\in\Sigma$, $\textbf{row}^{-\wedge}(\hat{\delta}(\textbf{row}(s),\sigma))=\delta(\textbf{row}^{-\wedge}(\textbf{row}(s)),\sigma)$.
Finally, we show that $\forall i\in\{k|k\in\mathbb{N}\And 1\leq k\leq|\Sigma^{O}|\},\textbf{row}^{-\wedge}:\hat{F}_{i}\rightarrow F_{i}$. Specifically, if $s\in S$ has $\textbf{row}(s)\in\hat{F}_{i}$, this means that $\textbf{row}(s)(\varepsilon)=\sigma_{i}$, and therefore $T(s)=T(s\cdot\varepsilon)=\sigma_{i}$. Furthermore, suppose $\textbf{row}^{-\wedge}(\textbf{row}(s))=\delta(q_{0},s)=q$, and therefore $\textbf{row}^{\wedge}(q)=\textbf{row}^{\wedge}(\delta(q_{0},s))\equiv\textbf{row}(s)$, which implies that $\textbf{row}^{\wedge}(q)(\varepsilon)=\textbf{row}(s)(\varepsilon)=T(s\cdot\varepsilon)=T(s)=\sigma_{i}$. This means that $q\in F_{i}$ and $\textbf{row}(s)\in\hat{F}_{i}$, and that each element of $\hat{F}_{i}$ bijectively maps to an element of $F_{i}$. ∎
Lemma 6.
Suppose $\langle S,E,T;\mathcal{C},\Gamma\rangle$ is a unified, closed, and consistent symbolic observation table. Let $n$ be the number of states in the hypothesis MakeHypothe-sis$(\langle S,E,T;\mathcal{C},\Gamma\rangle)$. Any Moore machine consistent with $\Gamma$ and $\mathcal{C}$ must have at least $n$ states.
Proof.
Let $h=\langle Q,\Sigma^{I},\Sigma^{O},q_{0},\delta,L\rangle$ be any Moore machine consistent with $T,\Lambda$, and $\mathcal{C}$, for the satisfying solution $\Lambda$ that was used to generate MakeHypothesis$(\langle S,E,T;\mathcal{C},\Gamma\rangle)$. Then, $\delta(q_{0},s)\in F_{i}$ if and only if $T(s)=\sigma_{i}$. Let $s_{1}$ and $s_{2}$ be distinct elements of $S$ such that $\textbf{{row}}(s_{1})\neq\textbf{{row}}(s_{2})$. This means there exists an $e\in E$ such that $T(s_{1}\cdot e)\neq T(s_{2}\cdot e)$. This means that if $T(s_{1}\cdot e)=\sigma_{i}$, then $T(s_{2}\cdot e)\neq\sigma_{i}$. Because $h$ is consistent with $T$, this means that $\delta(q_{0},s_{1}\cdot e)\in F_{i}$ and $\delta(q_{0},s_{2}\cdot e)\not\in F_{i}$. This means that $\delta(q_{0},s_{1}\cdot e)$ and $\delta(q_{0},s_{2}\cdot e)$ must be distinct since they must belong to different classes. Because $\delta(q_{0},s_{1}\cdot e)\neq\delta(q_{0},s_{2}\cdot e)$, and because the transition function must be consistent, then $\delta(q_{0},s_{1})$ and $\delta(q_{0},s_{2})$ must be distinct states: if $\delta(q_{0},s_{1})$ and $\delta(q_{0},s_{2})$ were the same state, then it is impossible to transition to two different states using the same sequence $e$. Since MakeHypothesis$(\langle S,E,T;\mathcal{C},\Gamma\rangle)$ contains $n$ states, there is at least one subset $P\subseteq S$ with $P=\{s_{1},...,s_{n}\}$ containing $n$ elements such that all the elements of $P^{\prime}=\{\textbf{{row}}(s_{1}),...,\textbf{{row}}(s_{n})\}$ are distinct from one another. Since every pair $\textbf{{row}}(s_{i})\neq\textbf{{row}}(s_{j})$ for $i\neq j$ in $P^{\prime}$, then it follows that each element of $\{\delta(q_{0},s_{1}),...,\delta(q_{0},s_{n})\}$ must be distinct to remain consistent with $T$. However, for a given $T$, there is a corresponding satisfying solution $\Lambda$ that specifies the values for $T$. This means that the mapping from $\Lambda$ to $T$ is one-to-one, and so for a given Moore machine to be consistent with a $(\Lambda,T)$ pair, the Moore machine must contain at least $n$ states. But since $\Lambda$ can be any satisfying solution to $\mathcal{C}$ and $\Gamma$, the above statement about Moore machines must be true for any $\Lambda$ that satisfies $\mathcal{C}$ and $\Gamma$. Thus, any Moore machine consistent with $\mathcal{C}$ and $\Gamma$ must have at least $n$ states. ∎
Theorem 1.
If the symbolic observation table $\langle S,E,T;\mathcal{C},\Gamma\rangle$ is unified, closed, and consistent, and $\mathcal{H}$ is the set of all hypotheses that can be returned from MakeHypothesis$(\langle S,E,T;\mathcal{C},\Gamma\rangle)$, then every hypothesis $h\in\mathcal{H}$ is consistent with constraints $\mathcal{C}$. Any other Moore machine consistent with $\mathcal{C}$, but not contained in $\mathcal{H}$, must have more states.
We first provide a sketch, followed by the proof.
Proof.
[…]
Proof.
Lemma 1 establishes that all hypotheses in $\mathcal{H}$ have equivalent states and equivalent transitions, which implies that every hypothesis in $\mathcal{H}$ has the same number of states. Lemmas 2 and 4 together establish that in a hypothesis, if a specific sequence $s$ is classified correctly, then all other sequences which start at the same initial state as $s$ and end in the same state as $s$ will be classified the same as $s$. Thus, if a state is classified correctly, then all sequences terminating at that state are classified correctly. Finally, Lemma 5 establishes that any Moore machine $d$ consistent with $\Lambda$ and $T$ from $\hat{h}=\textsc{MakeHypothesis}(\langle S,E,T;\mathcal{C},\Gamma\rangle)$ must be isomorphic to $\hat{h}$; otherwise $d$ must contain at least one more state than $\hat{h}$. This implies that $\hat{h}$ is the smallest Moore machine consistent with $\Lambda$ and $T$, and because $\Lambda$ is taken from all possible satisfying solutions to $\mathcal{C}$, then by Lemma 1, every hypothesis $h\in\mathcal{H}$ consistent with $\mathcal{C}$ has that same smallest number of states. Therefore, any other Moore machine not in $\mathcal{H}$, but still consistent with $\mathcal{C}$, must have more states. ∎
Theorem 1 establishes that the output of Algorithm 7 will be consistent with all the constraints in $\mathcal{C}$. In order to establish correctness and termination, we will present a few additional lemmas, and then the main result.
Lemma 7.
The number of unique constraints in $\mathcal{C}$ is finite, and is upper bounded by $\binom{|\Gamma|}{2}$.
Proof.
We will show that the number unique constraints in $\mathcal{C}$ is finite. In fact, the upper bound on the total number of constraints in $\mathcal{C}$ is a function of $|\Gamma|$, the total number of unique sequences recorded in the observation table. The total number of preference queries that are executed is $\binom{|\Gamma|}{2}$, which is the maximum number of unique constraints in $\mathcal{C}$.
This is because whenever a SymbolicFill is performed at some time $t$, there will be $N$ new, unique sequences in the table which have never been used in a preference query, and there are $O$ old sequences, representing sequences which have already been queried. At all times, $|\Gamma|=N+O$, and therefore, during a SymbolicFill, $\binom{N}{2}+NO$ preference queries are performed. Now, consider $r$ to represent the round number—the $r$th time that a SymbolicFill has been performed. Let $n_{r}$ represent the number of new, unique sequences in round $r$. Let $o_{r+1}$ represent all sequences in $\Gamma$ which have been queried over the past $r$ rounds; that is,
$o_{r+1}=\sum_{k=0}^{r}n_{k},$
where the initial conditions are $o_{0}=0$, and $n_{0}\geq 1$ is the initial size of $\Gamma$. Suppose that $R$ rounds have occurred. How many preference queries have occurred? Clearly, the total number of unique sequences after $R$ rounds is $n_{0}+n_{1}+\cdots+n_{R}=|\Gamma|$, which is the same as $o_{R}+n_{R}=|\Gamma|$. We can show that
$\binom{n_{0}+n_{1}+\cdots+n_{R}}{2}=\sum_{k=0}^{R}\binom{n_{r}}{2}+n_{r}o_{r}.$
Clearly, the RHS is justified, because in round $r$, $\binom{n_{r}}{2}+n_{r}o_{r}$ preference queries are performed. Therefore, we just need to show the LHS is equivalent to the RHS via some algebra:
$\binom{n_{0}+\cdots+n_{R}}{2}$ $=\binom{o_{R}+n_{R}}{2}$
$=\frac{(o_{R}+n_{R})(o_{R}+n_{R}-1)}{2}$
$=\frac{(o_{R}+n_{R})^{2}-(o_{R}+n_{R})}{2}$
$=\frac{o_{R}^{2}+n_{R}^{2}+2o_{R}n_{R}-o_{R}-n_{R}}{2}$
$=\frac{(o_{R}^{2}-o_{R})+(n_{R}^{2}-n_{R})+2o_{R}n_{R}}{2}$
$\binom{o_{R}+n_{R}}{2}$ $=\binom{o_{R}}{2}+\binom{n_{R}}{2}+o_{R}n_{R}$
$\binom{o_{R}+n_{R}}{2}$ $=\binom{o_{R-1}+n_{R-1}}{2}+\binom{n_{R}}{2}+o_{R}n_{R}$
$=\left[\binom{o_{R-2}+n_{R-2}}{2}+o_{R-1}n_{R-1}\right.$
$\left.+\binom{n_{R-1}}{2}\right]+\binom{n_{R}}{2}+o_{R}n_{R}$
Now, note
$\binom{o_{R-d}+n_{R-d}}{2}$ $=\binom{o_{R-d-1}+n_{R-d-1}}{2}$
$+\binom{n_{R-d}}{2}+o_{R-d}n_{R-d}$
for $10\leq d\leq R-$ 1 is an integer. Hence, by recursively splitting the above binomial term, we end up with the terms $o_{0}n_{0}+\cdots+o_{R}n_{R}$, $\binom{n_{0}}{2}+\cdots+\binom{n_{R}}{2}$, and $\binom{o_{0}}{2}$ in the expression, and in since $o_{0}=0$, we have:
$\binom{o_{R}+n_{R}}{2}$ $=\binom{o_{0}}{2}+\left[\binom{n_{0}}{2}+\cdots+\binom{n_{R}}{2}\right]$
$+(o_{0}n_{0}+\cdots+o_{R}n_{R})$
$=\binom{o_{0}}{2}+\sum_{k=0}^{R}\binom{n_{r}}{2}+o_{r}n_{r}$
$=\sum_{k=0}^{R}\binom{n_{r}}{2}+o_{r}n_{r}$
The number of unique sequences in the observation table is a function of the number of table closure and table consistency operations performed; but eventually, when the table becomes closed and consistent, a finite number of prefixes and suffixes will exist in the table. This implies there is only a finite number of unique possible sequences that can be created from the prefixes and suffixes in the table. The upper bound on the number of unique constraints is $\binom{|\Gamma|}{2}$. ∎
Lemma 8.
The number of representatives, $|\mathcal{R}|$, in a unified, closed, and consistent symbolic observation table $\langle S,E,T;\mathcal{C},\Gamma\rangle$ is finite, and is bounded by $1\leq|\mathcal{R}|\leq|\Sigma^{O}|$. Also, the number of unique inequalities in $\mathcal{C}$ is $\binom{|\mathcal{R}|}{2}$.
Proof.
Since the size of the output alphabet $\Sigma^{O}$ is finite, this means there is an upper bound of $|\Sigma^{O}|$ possible classification classes for all the sequences. If all sequences from $(S\cup(S\cdot\Sigma^{I}))\cdot E$ must be classified into $\Sigma^{O}$, then there are at most $|\Sigma^{O}|$ equivalence classes of sequences.
Because each entry of a unified, closed, and consistent symbolic observation table $\langle S,E,T;\mathcal{C},\Gamma\rangle$ is a variable, and that variable must have as its value one of the elements in $\Sigma^{O}$, the number of representatives $|\mathcal{R}|$ is bounded via $1\leq|\mathcal{R}|\leq|\Sigma^{O}|$. We define unique representatives to be representatives that are known have distinct values; in other words, the number of unique representatives at any given in time is the same as the number of equivalence classes over variables at that point in time; the unique representatives are the representatives of the equivalence classes of variables. Thus, if there are $|\mathcal{R}|$ unique representatives, and they satisfy a total ordering, then $\mathcal{C}$ will eventually contain a total of $\binom{|\mathcal{R}|}{2}$ unique inequalities. This total number of unique inequalities is populated in $\mathcal{C}$ via preference queries over pairs of unique sequences in the table. The quantity $|\Gamma|$ of unique sequences in the table is increased by prefix expansions (via closure tests and counterexamples), and suffix expansions (via consistency tests). Furthermore, during an equivalence query, whenever a counterexample $c$ is presented, $c$ and all its prefixes are added to the prefix set. ∎
We now define properties of a satisfying solution $\Lambda$ to the constraint set $C$ involving $|\mathcal{R}|$ representatives, and how many of those representatives have correct assignments. If $|\mathcal{R}|<V^{*}$ for some upper bound $V^{*}\leq|\Sigma^{O}|$, then we say that $\Lambda$ is incomplete. If $\Lambda$ contains at least 1 incorrect representative-value assignment, then we say $\Lambda$ is incorrect. If $\Lambda$ has assigned correct values to all $|\mathcal{R}|$ representatives, then we say $\Lambda$ is correct. If $\Lambda$ is both incomplete and correct, then we say $\Lambda$ is partially correct.
Lemma 9.
Whenever a counterexample $c$ is processed, either 00 or $1$ additional representative values becomes known.
Proof.
This is proven in the proof for Theorem 2. ∎
Theorem 2.
Suppose $\langle S,E,T;\mathcal{C},\Gamma\rangle$ is a unified, closed, and consistent symbolic observation table. Let $\hat{h}=\textsc{MakeHypothesis}(\langle S,E,T;\mathcal{C},\Gamma\rangle)$ be the hypothesis induced by $\Lambda$, a satisfying solution to $\mathcal{C}$. If the teacher returns a counterexample $c$ as the result of an equivalence query on $\hat{h}$, then at least one of the following statements about hypothesis $\hat{h}$ must be true: (a) $\hat{h}$ contains too few states, or (b) the satisfying solution $\Lambda$ inducing $\hat{h}$ is either incomplete or incorrect.
We first present a sketch, then a proof.
Proof.
(Sketch) This sketch applies to the above lemma, theorem, and below corollary about termination. Consider the sequence $\dots,h_{k-1},h_{k},\dots$ of hypotheses that Remap makes. For a given pair of consecutive hypotheses $(h_{k-1},h_{k})$, consider how the number of states $n$, and the number of known representative values $n_{\bullet}$ changes. Let $n^{*}$ be the number of states of the minimal Moore machine correctly classifying all sequences. Let $V^{*}\leq|\Sigma^{O}|$ be the upper bound on $|\mathcal{R}|$. Note that $0\leq n_{\bullet}\leq|\mathcal{R}|\leq V^{*}\leq|\Sigma^{O}|$ always holds. Through detailed case analysis on returned counterexamples, we can show that the change in $n_{\bullet}$, denoted by $\Delta n_{\bullet}$, must always be either 00 or $1$, and furthermore, if $\Delta n_{\bullet}=0$, then we must have $\Delta n\geq 1$. By the case analysis and tracking $n$ and $n_{\bullet}$, observe that if a counterexample $c$ is received from the teacher due to hypothesis $h$, then at least one of (a) $n<n^{*}$ or (b) $n_{\bullet}<V^{*}$ must be true. Since $\Delta n_{\bullet}$ and $\Delta n$ cannot simultaneously be 00, whenever a new hypothesis is made, progress must be made towards the upper bound of $(n^{*},V^{*})$. If the upper bound is reached, then the algorithm must terminate, since it is impossible to progress from the point $(n^{*},V^{*})$. ∎
Proof.
Our strategy for this proof is to show that the number of known representative values and number of states in the hypothesis will increase up to their upper limits, at which point the algorithm must terminate. Let $\mathbb{EC}[v]$ denote the set of variables in $\Gamma$ which are known to be equivalent to $v$ according to the equality constraints that were gathered so far in $\mathcal{C}$. Let $|\mathcal{R}|$ denote the number of representatives present in the unified, closed, and consistent observation table $\langle S,E,T;\mathcal{C},\Gamma\rangle$. That is, $|\mathcal{R}|$ is the number of elements in the set $\{\mathbb{EC}[\Gamma(s\cdot e)]|\forall s\cdot e\in(S\cup(S\cdot\Sigma))\cdot E\}$. Let $n_{\circ}$ represent the number of unknown representative values, and let $n_{\bullet}$ represent the number of known representative values. We have, at all times, the invariant $|\mathcal{R}|=n_{\circ}+n_{\bullet}$. Furthermore, at the very beginning of the algorithm, they have the initial values of $|\mathcal{R}|=1$, $n_{\circ}=|\mathcal{R}|$ and $n_{\bullet}=0$. Finally, let $V^{*}$ represent the ground truth number of classes that sequences can be classified into, according to the teacher. We will show that if the teacher returns a counterexample, then at least one of the following two scenarios must be true about the learner’s current hypothesis $\hat{h}=\textsc{MakeHypothesis}(\langle S,E,T;\mathcal{C},\Gamma\rangle)$: (A) $\hat{h}$ contains too few states, or (B) the satisfying solution $\Lambda$ inducing $\hat{h}$ is incorrect. We enumerate cases by describing the pre-conditions and post-conditions when executing the equivalence query.
Case 1.
Prior to the equivalence query, $c\in(S\cup(S\cdot\Sigma^{I}))\cdot E$. This means there exists $s_{1}\in(S\cup(S\cdot\Sigma^{I}))$ and $e_{1}\in E$ such that $c=s_{1}\cdot e_{1}$ and the variable located at $T(c)$ was assigned an incorrect value $w$. This also means that for every $s\cdot e\in(S\cup(S\cdot\Sigma))\cdot E$ such that $\Gamma[s\cdot e]\in\mathbb{EC}[T(c)]$, it follows that each $T(s\cdot e)$ is also the same incorrect $w$. However, after the equivalence query, the teacher returns the counterexample $c$ with feedback $f(c)=r$, where $r\neq~{}w$. Then, for each variable $v\in\mathbb{EC}[T(c)]$, the learner will assign $v$ the value $r$. If $c\in S$, then all of its prefixes are already in $S$ because $S$ is prefixed closed. If $c\in S\cdot\Sigma$, then all of its prefixes are already in $S$. If $c$ is not in $S\cdot\Sigma$, then the prefixes of $c$ which are not in $S$ will be added to $S$. In this latter case, there is a possibility that the number of unique representatives $|\mathcal{R}|$ will be increased by $k$, where $0\leq k\leq V^{*}-|\mathcal{R}|$. However, the number of known variables will always increase by exactly 1. Thus, under this case, we have the following transformation: $(|\mathcal{R}|,n_{\circ},n_{\bullet})\longrightarrow(|\mathcal{R}|+k,n_{\circ}-1+k,n_{\bullet}+1)$, where $0\leq k\leq V^{*}-|\mathcal{R}|$.
Case 2.
Prior to the equivalence query, $c\not\in(S\cup(S\cdot\Sigma^{I}))\cdot E$, which means $T(c)$ does not exist in the symbolic observation table. If $c$ was returned as a counterexample, then the learner incorrectly hypothesized that $\hat{\delta}(\hat{q}_{0},c)\in\hat{F}_{i}$. The post-conditions for this case can be broken into three possibilities:
  1. (a)
    There is some $s\cdot e\in(S\cup(S\cdot\Sigma^{I}))\cdot E$ for which $\mathbb{EC}\left[\Gamma[s\cdot e]\right]\overset{\mathrm{val}}{=}r$ according to $\Lambda$, and $\left|\mathbb{EC}[\Gamma[s\cdot e]]\right|\geq 1$. The new variable $T(c)$ is added to $\mathbb{EC}[\Gamma[s\cdot e]]$, and subsequently, $\mathbb{EC}[T(c)]=\mathbb{EC}[\Gamma[s\cdot e]]$ and $|\mathbb{EC}[T(c)]|>1$, and all the variables in that class have a value $r$. While $c$ itself does not induce $|\mathcal{R}|$ to increase, it is possible that its prefixes might induce an increase in $|\mathcal{R}|$. So while the number of unknown representative values might increase, the number of known representative values stays constant. Therefore, we have the following transformation: $(|\mathcal{R}|,n_{\circ},n_{\bullet})\longrightarrow(|\mathcal{R}|+k,n_{\circ}+k,n_{\bullet})$, where $0\leq k\leq V^{*}-|\mathcal{R}|$.
  2. (b)
    There is some $s\cdot e\in(S\cup(S\cdot\Sigma^{I}))\cdot E$ for which $\mathbb{EC}\left[\Gamma[s\cdot e]\right]\overset{\mathrm{val}}{\neq}r$ according to $\Lambda$ and $\left|\mathbb{EC}[\Gamma[s\cdot e]]\right|\geq 1$. The new variable $T(c)$ is added to $\mathbb{EC}[\Gamma[s\cdot e]]$, and subsequently, $\mathbb{EC}[T(c)]=\mathbb{EC}[\Gamma[s\cdot e]]$ and $|\mathbb{EC}[T(c)]|>1$, and $\mathbb{EC}\left[T(c)\right]\overset{\mathrm{val}}{=}\mathbb{EC}\left[\Gamma[s\cdot e]\right]\overset{\mathrm{val}}{=}r$ Again, the number of unknown representative values might increase, but this time, the number of known representative values increases by 1. Therefore, we have the following transformation: $(|\mathcal{R}|,n_{\circ},n_{\bullet})\longrightarrow(|\mathcal{R}|+k,n_{\circ}-1+k,n_{\bullet}+1)$, where $0\leq k\leq V^{*}-|\mathcal{R}|$.
  3. (c)
    The variable $T(c)$ is not added to any pre-existing equivalence class; therefore it is added to its own equivalence class, and its value is known to be $r$. Thus, we have the following transformation: $(|\mathcal{R}|,n_{\circ},n_{\bullet})\longrightarrow(|\mathcal{R}|+1+k,n_{\circ}+k,n_{\bullet}+1)$, where $10\leq k\leq V^{*}-|\mathcal{R}|-$ 1.
We note that cases (1), (2b), and (2c) can occur a finite number of times. This is because $n_{\bullet}$ increases by exactly $1$ each time one of those cases occurs. The maximum value $n_{\bullet}$ can take on is $|\mathcal{R}|$; the upper bound on $|\mathcal{R}|$ is $V^{*}$, and the upper bound on $V^{*}$ is $|\Sigma^{O}|$. Furthermore, cases (1), (2b), and (2c) cannot occur if $n_{\bullet}=V^{*}$. Thus, if case (1), (2b), or (2c) occur, they increase the number of known variables by 1, and they will cease to occur once $n_{\bullet}=V^{*}$. These cases fall under the umbrella of an incorrect satisfying solution $\Lambda$.
Clearly, case (1) falls under Scenario (B), an incorrect satisfying solution $\Lambda$, since the value of $T(c)$ for a known $c$ in the table is incorrect. Case (2b) also falls under Scenario (B), because $\hat{\delta}(\hat{q}_{0},c)=\hat{\delta}(\hat{q}_{0},s^{\prime})$ for some $s^{\prime}\in S$ was incorrectly classified, implying that the value of $T(s^{\prime})$ was incorrect.
For cases (2a) and (2c), it is not necessarily true that the satisfying solution $\Lambda$ used to generate $\hat{h}$ was incorrect, with respect to the current number of representatives $|\mathcal{R}|$. It is possible that $\Lambda$ was indeed incorrect, where at least one of the $|\mathcal{R}|$ representatives was assigned the incorrect value, in which case Scenario (B) is true. It is also possible that $\Lambda$ was incomplete, but partially correct, where all existing known representatives had correct values, but where $|\mathcal{R}|<V^{*}$. If a hypothesis $\hat{h}$ is partially correct, then $\mathcal{C}$ contains only $|\mathcal{R}|$ representatives, with $|\mathcal{R}|<V^{*}$, $n_{\bullet}=|\mathcal{R}|$, and $n_{\circ}=0$. This means that $\hat{h}=\textsc{MakeHypothesis}(\langle S,E,T;\mathcal{C},\Gamma\rangle)$ is the only partially correct hypothesis the algorithm could have generated under these conditions (any others must be incorrect), and is therefore unique; the size of $\mathcal{H}$ must be 1. Then, by Theorem 1, any other Moore machine consistent with $\mathcal{C}$, but not equivalent to $\hat{h}$ must contain more states; hence Scenario (A) is true. Thus, in our case work, we have shown that Scenario (A) or Scenario (B) holds. ∎
Corollary 1 (Termination).
Remap must terminate when the number of states and number of known representative values in a concrete hypothesis reach their respective upper bounds.
Proof.
Since cases (1-2c) cover all the cases, and in each case at least one of the scenarios, Scenario (A) and Scenario (B), is true, then for each pair of consecutive hypotheses $\hat{h}_{j-1}$ and $\hat{h}_{j}$ generated by consecutive equivalence queries in the algorithm, one of the following is true: (a) the latter hypothesis $\hat{h}_{j}$ contains at least one more state than the prior hypothesis $\hat{h}_{j-1}$, (b) the latter hypothesis $\hat{h}_{j}$ contains at least one more known variable value compared to the prior hypothesis $\hat{h}_{j-1}$, or (c) both (a) and (b) are true. Thus, in a sequence of hypotheses generated by equivalence queries, both the number of states and number of known representatives increase monotonically. There can be at most $V^{*}$ discoveries. If the $n^{*}$ is the number of states in the minimum Moore machine which correctly classifies all sequences, then the number of states $n$ in each hypothesis will increase monotonically to $n^{*}$. Then, clearly, the algorithm must terminate when $n=n^{*}$ and $|\mathcal{R}|=V^{*}$. ∎
Corollary 1 indicates that Remap must terminate, since every hypothesis made makes progress towards the upper bound.
Theorem 3 (Query Complexity).
If $n$ is the number of states of the minimal automaton isomorphic to the target automaton to be learned, and $m$ is the maximum length of any counterexample sequence that the teacher returns, then (a) the upper bound on the number of equivalence queries that Remap executes is $n+|\Sigma^{O}|-1$, and (b) the preference query complexity is $\mathcal{O}(mn^{2}\ln(mn^{2}))$, which is polynomial in the number of unique sequences that the learner performs queries on.
Proof.
Based on Theorem 2, we know that the maximum number of equivalence queries is the taxi distance from the point $(1,0)$ to $(n,|\Sigma^{O}|)$, which is $n+|\Sigma^{O}|-1$. From counterexample processing, we know there will be at most $m(n+|\Sigma^{O}|-1)$ sequences added to the prefix set $S$, since a counterexample $c$ of length $m$ results in at most $m$ sequences added to the prefix set $S$. The maximum number of times the table can be found inconsistent is at most $n-1$ times, since there can be at most $n$ states, and the learner starts with $1$ state. Whenever a sequence is added to the suffix set $E$, the maximum length of sequences in $E$ increases by at most $1$, implying the maximum sequence length in $E$ is $n-1$. Similarly, closure operations can be performed at most $n-1$ times, so the total number of sequences in $E$ is at most $n$; the maximum number of sequences in $S$ is $n+m(n+|\Sigma^{O}|-1)$. The maximum number of unique sequences queried in the table is the maximum cardinality of $(S\cup S\cdot\Sigma^{I})\cdot E$, which is
$(n+m(n+|\Sigma^{O}|-1))(1+|\Sigma^{I}|)n=\mathcal{O}(mn^{2}).$
Therefore, the preference query complexity of Remap is $\mathcal{O}(mn^{2}\ln(mn^{2}))$ due to randomized quicksort. ∎
Next, we show that Remap probably approximately correctly identifies the minimal automaton isomorphic to the target.
Definition 6 (Probably Approximately Correct Identification).
Given an arbitrary Moore machine $M=\langle Q,q_{0},\Sigma^{I},\Sigma^{O},\delta,L\rangle$, let the regular language classification function $f:(\Sigma^{I})^{*}\rightarrow\Sigma^{O}$ be represented by $f(s)=L(\delta(q_{0},s))$ for all $s\in(\Sigma^{I})^{*}$. Let $\mathcal{D}$ be an any probability distribution over $(\Sigma^{I})^{*}$. An algorithm $\mathcal{A}$ probably approximately correctly identifies $f$ if and only if for any choice of $10<\epsilon<$ 1 and $10<d<$ 1, $\mathcal{A}$ always terminates and outputs an $\epsilon$-approximate sequence classifier $\hat{f}:(\Sigma^{I})^{*}\rightarrow\Sigma^{O}$, such that with probability at least $1-d$, the probability of misclassification is $P(\hat{f}(s)\neq f(s))\leq\epsilon$ when $s$ is drawn according to the distribution $\mathcal{D}$.
Theorem 4.
Remap achieves probably approximately correct identification of any Moore machine when the teacher $\mathcal{T}$ uses sampling-based equivalence queries.
We first present a proof sketch, then the full proof.
Proof.
(Sketch) The probability $1-\epsilon_{k}$ of a sequence sampled from an arbitrary distribution $\mathcal{D}$ over $(\Sigma^{I})^{*}$ depends on the distribution and the intersections of sets of sequences of the teacher and the learner’s $k$th hypothesis with the same classification values. The probability that the $k$th hypothesis misclassifies a sequence is $\epsilon_{k}$. If the teacher samples $m_{k}$ samples for the $k$th equivalence query, then an upper bound can be established for the case when $\epsilon_{k}\leq\epsilon$ for a given $\epsilon$. Since we know Remap executes at most $n+|\Sigma^{O}|-1$ equivalence queries, one can upper bound the probability that Remap terminates with an error by summing all probabilities of events that the teacher does not detect an error in at most $n+|\Sigma^{O}|-1$ equivalence queries. An exponential decaying upper bound can be found, and a lower bound for $m_{k}$ can be found in terms of $\epsilon,d,$ and $k$. ∎
Proof.
Suppose the teacher $\mathcal{T}$ has Moore machine $\langle Q,q_{0},\Sigma^{I},\Sigma^{O},\delta,L\rangle$ representing $f$ which classifies all sequences in the set $(\Sigma^{I})^{*}$ into $|\Sigma^{O}|$ disjoint sets $\{L_{c}|c\in\Sigma^{O}\}$ such that for all $c\in\Sigma^{O},s\in L_{c},f(s)=L(\delta(q_{0},s))=c$, we have:
$(\Sigma^{I})^{*}=\bigcup_{c\in\Sigma^{O}}L_{c}.$
Similarly, suppose the learner has proposed a Moore machine $\langle\hat{Q},\hat{q}_{0},\Sigma^{I},\Sigma^{O},\hat{\delta},\hat{L}\rangle$ representing the $k$th hypothesis $\hat{f}_{k}$ which classifies all sequences in the set $(\Sigma^{I})^{*}$ into $|C^{k}_{L}|$ disjoint subsets $\{L^{\prime k}_{c}|c\in C^{k}_{L}\}$ where $C^{k}_{L}\subseteq\Sigma^{O}$ such that for all $c\in C^{k}_{L}$ and for all $s\in L^{\prime k}_{c},\hat{f}(s)=\hat{L}(\hat{\delta}(\hat{q}_{0},s))=c$ we have:
$(\Sigma^{I})^{*}=\bigcup_{c\in C^{k}_{L}}L^{\prime k}_{c}$
If $\mathcal{D}$ is a distribution over $(\Sigma^{I})^{*}$, and if $S\sim\mathcal{D}$ is a random variable representing the sequence $s$ drawn according to $\mathcal{D}$, and if the set of intersections $I_{k}$ is defined by
$I_{k}=\displaystyle\bigcup_{c\in C^{k}_{L}\cap\Sigma^{O}}L_{c}\cap L^{\prime k}_{c}$
then the probability that the $k$th hypothesis $\hat{f}_{k}$ classifies $s$ correctly is
$P(\hat{f}_{k}(S)=f(S))=\displaystyle\sum_{s\in I_{k}}p_{S}(s)$
and therefore the probability $\epsilon_{k}$ that $\hat{f}_{k}(s)\neq f(s)$ is
$\epsilon_{k}=P(\hat{f}_{k}(S)\neq f(S))=1-\displaystyle\sum_{s\in I}p_{S}(s).$
Suppose for the $k$th equivalence query, the teacher $\mathcal{T}$ samples $m_{k}$ sequences i.i.d. according to the distribution $\mathcal{D}$ over $(\Sigma^{I})^{*}$. The probability $\mathcal{T}$ accepts $\hat{f}_{k}$ because it detects no misclassification for any of the $m_{k}$ samples (represented by the random variables $S_{1},\cdots,S_{m_{k}}\sim\mathcal{D}$) is $p_{k}$, given by
$p_{k}=P(\sum_{i=1}^{m_{k}}(\hat{f}_{k}(S_{i})-f(S_{i}))=0)=(1-\epsilon_{k})^{m_{k}},$
so if $\epsilon_{k}\leq\epsilon$ for a chosen $\epsilon$, then
$P(\sum_{i=1}^{m_{k}}(\hat{f}_{k}(S_{i})-f(S_{i}))=0|\epsilon_{k}\leq\epsilon)\geq(1-\epsilon)^{m_{k}}$
and if $\epsilon_{k}\geq\epsilon$, then
$P(\sum_{i=1}^{m_{k}}(\hat{f}_{k}(S_{i})-f(S_{i}))=0|\epsilon_{k}\geq\epsilon)\leq(1-\epsilon)^{m_{k}}.$
We know that Remap will execute at most $n+|\Sigma^{O}|-1$ equivalence queries, so the probability that Remap terminates with an error in the set $\{\epsilon_{1},\cdots,\epsilon_{n+|\Sigma^{O}|-1}\}$ is given by
$p_{1}+(1-p_{1})p_{2}+\cdots+p_{n+|\Sigma^{O}|-1}\prod_{k=1}^{n+|\Sigma^{O}|-2}(1-p_{k})$
$=\sum_{k=1}^{n+|\Sigma^{O}|-1}p_{k}\prod_{i=0}^{k-1}(1-p_{i})$
where $p_{0}=0$. Each term in the summation on the right hand side is less than or equal to $p_{k}$, so
$\sum_{k=1}^{n+|\Sigma^{O}|-1}p_{k}\prod_{i=0}^{k-1}(1-p_{i})\leq\sum_{k=1}^{n+|\Sigma^{O}|-1}p_{k}.$
Furthermore, we know that $\epsilon_{k}$ monotonically decreases as $k$ increases because $|C_{L}^{k}\cap\Sigma^{O}|$ monotonically increases. Considering the progress transformation from Theorem 2, if $|C_{L}^{k}\cap\Sigma^{O}|$ monotonically increases, then this means the cardinality of the set $\{{L^{\prime}}_{c}^{k}|\forall c\in C_{L}^{k}\}$ monotonically increases. If any sequence in any of the intersections $L_{c}\cap{L^{\prime}}_{c}^{k}$ is misclassified, then the feedback from the counterexample will move the misclassified sequence to the correct class, which means the error decreases monotonically. Specifically, by Theorem 2, the following events can occur: (1) the value of an existing state (or a set of existing equivalently valued states) will obtain their correct values; (2) at least 1 new states will be created, either increasing the size of one of the ${L^{\prime}}_{c}^{k}$; or (3) increasing the cardinality of $|C_{L}^{k}\cap\Sigma^{O}|$ and creating an additional ${L^{\prime}}_{c}^{k}$ set (corresponds with at least one new states and one new value). Other non-counterexample sequences which were previously misclassified either stay misclassified (perhaps with a different, but still incorrect value), or they will become classified correctly. A sequence which was already known to be classified correctly will not become misclassified. Thus, all three possibilities monotonically decrease the error. This means $\epsilon_{j}\geq\epsilon_{j+1}$, so the terminal hypothesis will have an error of at most the error of the previous hypothesis. If we choose a desired error $\epsilon$, then the probability that the error of the terminal hypothesis is greater than $\epsilon$ is at most
$\sum_{k=1}^{n+|\Sigma^{O}|-1}p_{k}\leq\sum_{k=1}^{n+|\Sigma^{O}|-1}(1-\epsilon)^{m_{k}}\leq\sum_{k=1}^{n+|\Sigma^{O}|-1}e^{-\epsilon m_{k}}$
since
$P(\sum_{i=1}^{m_{k}}(\hat{f}_{k}(S_{i})-f(S_{i}))=0|\epsilon_{k}\geq\epsilon)\leq(1-\epsilon)^{m_{k}},$
and since $1+x\leq e^{x}$ for all real $x$. If we take each term in the rightmost summation and require it to be at most $\frac{d}{2^{k}}$ for some chosen value of $10<d<$ 1, then we have a lower bound on the number of samples for the $k$th equivalence query for $k=1,\cdots,(n+|\Sigma^{O}|-1)$
$m_{k}\geq\frac{1}{\epsilon}(\ln\frac{1}{d}+k\ln 2)$
and
$\sum_{k=1}^{n+|\Sigma^{O}|-1}e^{-\epsilon m_{k}}\leq\sum_{k=1}^{n+|\Sigma^{O}|-1}\frac{d}{2^{k}}\leq\sum_{k=1}^{\infty}\frac{d}{2^{k}}\leq d$
which implies shows that the probability that Remap terminates with a hypothesis with error at least $\epsilon$ is at most $d$. ∎
Theorem 5.
To achieve PAC-identification under Remap, if the teacher $\mathcal{T}$ has chosen parameters $\epsilon$ and $d$, and if $f$ can be represented by a minimal Moore machine with $n$ states and $|\Sigma^{O}|$ classes, then $\mathcal{T}$ needs to sample at least
$\mathcal{O}(n+|\Sigma^{O}|+\frac{1}{\epsilon}((n+|\Sigma^{O}|)\ln\frac{1}{d}+(n+|\Sigma^{O}|)^{2}))$
sequences i.i.d. from $\mathcal{D}$ over the entire run of Remap.
Proof.
Since for the $k$th equivalence query, the teacher must sample at least $m_{k}\geq\left\lceil\frac{1}{\epsilon}(\ln\frac{1}{d}+k\ln 2)\right\rceil$ sequences in order to achieve PAC-identification, if the total number of samples is to be minimized while still achieving PAC-identification, then the teacher can just sample a total of
$\sum_{k=1}^{n+|\Sigma^{O}|-1}\left[\frac{1}{\epsilon}(\ln\frac{1}{d}+k\ln 2)+1\right]$
$=$ $n+|\Sigma^{O}|-1$
$+\frac{1}{\epsilon}\left[(\ln\frac{1}{d})(n+|\Sigma^{O}|-1)+\ln 2\sum_{k=1}^{n+|\Sigma^{O}|-1}k\right]$
$=$ $\mathcal{O}\left((n+|\Sigma^{O}|)+\frac{1}{\epsilon}((n+|\Sigma^{O}|)\ln\frac{1}{d}+(n+|\Sigma^{O}|)^{2})\right)$
Miscellaneous Experiments
In addition to learning reward machines, we also evaluated Remap on sequence classification problems where the classification model can be represented as a Moore machine. Since reward machines are Mealy machines, and Moore machines and Mealy machines are equivalent, each of these Moore machines can be converted into a reward machine. We specified the reference Moore machines using regular expressions. The regular expressions we used to construct the Moore machines were: $a^{*}b$ represents sequences of zero or more $a$’s followed by a single $b$; $b^{*}a$ represents sequences of zero or more $b$’s followed by a single $a$; $(a|b)^{*}$ represents sequences with zero or more elements, where each element is either $a$ or $b$; $(ab)^{*}$ represents sequences of zero or of even lengths, where every $a$ is immediately followed by a $b$, and non-zero length sequences start with an $a$; $(ba^{*})$ represents sequences of zero or of even lengths, where every $b$ is immediately followed by a $a$, and non-zero length sequences start with a $b$.
These components were OR’ed together. If $M_{i}$ represents a regular expression component, and
$M=\displaystyle\left|{}_{i=1}^{N}M_{i}\right.=M_{1}|M_{2}|\cdots|M_{N}$
represents $N$ distinct regular expression components OR’ed together, then the classifier $f$ classifies a sequence $s$ in the following manner: $f(s)=k$ if $s\in M_{k}$, and is 00 otherwise. If $s$ happens to be contained in multiple $M_{k_{1}}$,$M_{k_{2}},...$, then the lowest matching $k$ value prevails. This structure implies the size of the output alphabet is $1+N$.
We measured accuracy, number of preference queries, number of equivalence queries, and number of unique sequences as functions of samples per equivalence query (Figure 9) and Figure 10. We also measured these attributes as functions of alphabet size (Figure 11). Finally, we also created some termination phase diagrams (Figure 12).
{tblr}
colspec = X[c]X[c,h]X[c,h]X[c,h]X[c,h], stretch = 0, rowsep = 6pt, Regex & Accuracy Num of EQ Num of PQ Num of Seq
$(a^{*}b)$ Refer to caption Refer to caption Refer to caption Refer to caption
$(a^{*}b)|(b^{*}a)$ Refer to caption Refer to caption Refer to caption Refer to caption
$(a^{*}b)|(b^{*}a)|(a|b)^{*}$ Refer to caption Refer to caption Refer to caption Refer to caption
$(a^{*}b)|(b^{*}a)|(ab)^{*}$ Refer to caption Refer to caption Refer to caption Refer to caption
$(a^{*}b)|(b^{*}a)|(ab)^{*}|(ba)^{*}$ Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Violin Plots of Terminal Hypothesis Accuracy, Number of Equivalence Queries executed, Number of Preference Queries executed, and Number of Unique Sequences Tested, as a function of Maximum Number of Random Tests performed per Equivalence Query, for learning a variety of regexes using a two-letter alphabet. During an Equivalence Query, the teacher samples $N$ random sequences, with the length of each sequence drawn iid from a geometric distribution with termination probability $0.2$. Given a sequence test length, each element of the sequence is chosen iid from a uniform distribution over the alphabet. To compute the terminal hypothesis accuracy, the terminal hypothesis is tested on $200$ random sequences per regex class. A total of $2000$ trials were conducted per $N$ value; the plots show the resulting empirical distributions of those trials.
Figure 10: Plots of the same data as in Figure 9. Mean values are represented by thick lines, median values are represented by thin lines, and values lying between the $20^{th}$ and $80^{th}$ percentiles are shown as shaded.
Figure 11: Plots accuracy, number of unique sequences tested, number of equivalence queries performed, and number of preference queries made, as a function of alphabet size. These plots are set at 200 samples per EQ. Mean values are represented by thick lines, median values are represented by thin lines, and values lying between the $20^{th}$ and $80^{th}$ percentiles are shown as shaded.
Figure 12: Termination plots illustrating that number of state and number of known variables increase monotonically. The blue circles represent closure operations, orange squares represent consistency operations, and green Xs represent equivalence queries. The number of states in the automata from left to right is 3, 5, 8, 11, 14. We observe that for termination to occur, the number of states must be correct. The number of explicitly known values does not need to match the number of classes to terminate, however, because the learner can simply “guess” the correct values for early termination.